### v2

# C-Learner: Constrained Learning for Causal Inference and Semiparametric Statistics

Abstract
Causal estimation (e.g. of the average treatment effect) requires estimating complex nuisance parameters (e.g. outcome models). To adjust for errors in nuisance parameter estimation, we present a novel correction method that solves for the best plug-in estimator under the constraint that the first-order error of the estimator with respect to the nuisance parameter estimate is zero. Our constrained learning framework provides a unifying perspective to prominent first-order correction approaches including one-step estimation (a.k.a. augmented inverse probability weighting) and targeting (a.k.a. targeted maximum likelihood estimation). Our semiparametric inference approach, which we call the “C-Learner”, can be implemented with modern machine learning methods such as neural networks and tree ensembles, and enjoys standard guarantees like semiparametric efficiency and double robustness. Empirically, we demonstrate our approach on several datasets, including those with text features that require fine-tuning language models. We observe the C-Learner matches or outperforms other asymptotically optimal estimators, with better performance in settings with less estimated overlap.
1 Introduction
Causal inference is the bedrock of scientific decision-making. To reliably estimate causal effects, a modeler often needs to fit high-dimensional and complex nuisance parameters. For example, optimal estimation of the average treatment effect (ATE) requires modeling outcomes and/or treatment assignments as a function of baseline covariates, even for a randomized controlled trial [21]. Modern machine learning models provide substantial flexibility in fitting nuisance parameters, but they optimize predictive accuracy and are agnostic to the downstream estimation task. As a result, a naive plug-in estimator that entirely trusts the fitted ML model is suboptimal and sensitive to errors in the ML model [15, 19, 10].
To improve upon the naive plug-in estimator, the semiparametric statistics literature analyzes the sensitivity of the causal estimand with respect to the nuisance parameter by taking a first-order distributional Taylor expansion [4, 26, 27, 22, 24]. There are two prominent approaches to correcting the first-order error term due to nuisance parameter estimation. One-step estimation corrects the plug-in estimator by subtracting (an estimate of) the first-order error term, which can be estimated using the pathwise derivative of the estimand (a.k.a. efficient influence function). On the other hand, targeting uses the plug-in estimator but adjusts the nuisance parameter estimate by a pathwise derivative-based (“clever”) covariate term designed to set (an estimate of) the first-order estimation error to zero [40]. In the case of estimating the ATE, the two approaches respectively give rise to the well-known augmented inverse probability weighting (AIPW) estimator [7, 33, 31] and the targeted maximum likelihood estimator (TMLE) [40] which are both semiparametrically efficient and doubly robust.
We present a new approach to correcting the first-order estimation error in causal inference and semiparametric statistics, which we make concrete in the case of estimating the ATE. Our approach reframes the problem of constructing an asymptotically optimal ATE estimator as a constrained optimization problem. Consider binary treatments (actions) $A$, covariates $X$, and potential outcomes $Y(1),Y(0)$ under treatment ($A=1$) and control ($A=0$), respectively. Under standard identification assumptions for the ATE, letting $(X,A,Y)\sim P$ for some distribution $P$, and also assuming $Y(0):=0$ for simpler exposition, the ATE depends on the outcome model $\mu(\cdot)$:
$\mbox{ATE}=\mathbb{E}_{P}[Y(1)-Y(0)]=\mathbb{E}_{P}[Y(1)]=\mathbb{E}_{P}[\mu(X)]~{}~{}~{}\mbox{where}~{}~{}~{}\mu(X):=\mathbb{E}_{P}[Y\mid A=1,X].$
Instead of fitting $\mu(\cdot)$ to minimize predictive error, we explicitly take into account the downstream causal estimation task by minimizing predictive error subject to the constraint that the first-order estimation error must be zero
$\widehat{\mu}^{C}\in\mbox{argmin}_{\tilde{\mu}\in\mathcal{F}}\Bigg{\{}\mbox{PredError}(\tilde{\mu}):\mbox{First-order error of plug-in estimator from }\tilde{\mu}\text{ is }0\Bigg{\}}$ (1)
where the minimization is over the class of outcome models $\tilde{\mu}(\cdot)$ over the model class $\mathcal{F}$.
The resulting plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}^{C}(X_{i})$ based on the constrained learning approach (1) enjoys the usual fruits of first-order correction, such as semiparametric efficiency and double robustness. Our theory provides a unifying perspective for standard first-order correction methods like one-step estimation and targeting. These existing methods can be thought of as solving the optimization problem (1) with a restrictive model class given by augmenting an fitted nuisance estimator along a specific direction. For estimating the ATE, a version of one-step estimation uses an existing outcome model $\widehat{\mu}(X)$ plus an additive constant term :=$\mathcal{F}:=\{\widehat{\mu}(\cdot)+c:c\in\mathbb{R}\}$, and a version of targeting uses a pathwise derivative-based covariate term :=$\mathcal{F}:=\{\widehat{\mu}(\cdot)+c\frac{A}{\widehat{\pi}(\cdot)}:c\in\mathbb{R}\}$ where $A$ is treatment and $\widehat{\pi}$ is a fitted propensity score model (treatment model). We elaborate further in Section 3.1.
More importantly, the constrained learning perspective (1) gives rise to a natural and novel estimation approach, which we call the “C-Learner”. Instead of adjusting an existing nuisance parameter estimate along a pre-specified direction, we directly train the nuisance parameter to minimize prediction error, subject to the constraint that the first-order estimation error must be zero. Unlike one-step estimation and targeting, our approach utilizes the entire class of ML models to make the first-order correction. Intuitively, we expect the C-Learner to enjoy practical finite-sample advantages over one-step estimation and targeting by virtue of directly optimizing the plug-in model over this originally chosen model class $\mathcal{F}$. In practice, we observe the plug-in estimator using outcome models fitted under the constrained framework to exhibit numerical stability.
Since the constrained learning problem (1) is a first-order correction method, the resulting plug-in estimator satisfies standard asymptotic optimality guarantees such as semiparametric efficiency and double robustness. In Section 6, we identify conditions under which a constrained learning estimator enjoys these results, and we verify that one-step estimation and targeting methods also satisfy these necessary conditions. While all first-order correction methods enjoy standard asymptotic optimality guarantees, several authors have noted that there is a salient gap between the asymptotic and finite-sample performance of different estimation approaches [23, 8]. This leaves room for new methodological development (including ours) and necessitates a rigorous and thorough empirical comparison.
We demonstrate the versatility of the C-Learner by providing implementations in which the outcome models are linear models, gradient-boosted regression trees, or neural networks (Seection 4). We study a range of datasets where different outcome models are appropriate: standard tabular settings where linear models and gradient-boosted trees excel, and a large-scale setting involving text features that require finetuning language models for causal estimation. Empirically, we find the C-Learner outperforms or matches other asymptotically optimal estimation methods (Section 5). In particular, in our experiment settings, we observe that the C-Learner has better estimation error in challenging settings where there are covariate regions with little estimated overlap between treatment and control groups.
Machine learning approaches for nuisance estimation have received much attention [1, 2, 3, 29, 18, 44]. These approaches can be naturally combined with first-order correction methods like one-step estimation and targeting, and recent advancements including DragonNet [38] and RieszNet [12] explicitly consider these ideas during model training through novel objectives and model classes. Like other first-order correction methods, the C-Learner can be combined with various nuisance model estimation approaches; for example, in Appendix C.3.2, we demonstrate how C-Learner can effectively use Riesz representers learned by RieszNet [12]. Our constrained optimization view reframes the construction of an asymptotically optimal estimator as a constrained optimization problem, and provides new ways to achieve asymptotic optimality in practice.
2 Background
Although C-Learner is applicable to more general estimands (Section A), we illustrate our ideas by studying the average treatment effect (ATE).
2.1 Average Treatment Effect and Missing Outcomes
[…]
To simplify our exposition, we assume $Y(0):=0$ throughout so that we are estimating the mean of a censored outcome (censored when $A=0$). This setting is also known as mean missing outcome [24], which is the focus of some of our experiments. In this case, we can write the ATE as a functional of the joint measure $P$, as
$\psi(P):=P[Y(1)]=P\left[P[Y\mid A=1,X]\right].$ (2)
The outcome model $\mu(X):=P[Y\mid A=1,X]$ (shorthand for $\mu(A,X):=P[Y\mid A,X]$ with $\mu(0,X):=0$) and propensity (treatment) model $\pi(X):=P(A=1\mid X)$ are key nuisance parameters; notably, they are high-dimensional in contrast to the single-dimensional ATE (2). Note that we can write $\psi(P)=P[\mu(1,X)-\mu(0,X)]=P[\mu(1,X)]=P[\mu(X)]$ in this setting. The corresponding nuisance estimators $\widehat{\mu}(X)$ and $\widehat{\pi}(X)$ can be implemented as ML models that are trained on held-out data.
2.2 First-Order Correction and Asymptotic Optimality
By analyzing the error of blindly trusting an ML-based estimate of the nuisance parameters to estimate the ATE (2), we can develop better estimation approaches. We sketch intuition here and leave a more rigorous treatment to Section 6. We begin by noting that the joint distribution over the observed data can be decomposed as $P=P_{Y,A|X}\times P_{X}$. Since the marginal $P_{X}$ can be simply estimated with an empirical distribution (plug-in), we focus on the error induced by approximating $P_{Y,A|X}$ using an estimate $\widehat{P}_{Y,A|X}$.
Recalling the statistical functional $\psi(P)$ (2), the error of a naive plug-in estimator, assuming access to the true $P_{X}$, is given by
$P_{X}[\widehat{\mu}(X)]-P_{X}[\mu(X)]$ $=P_{X}[\widehat{P}[Y\mid A=1,X]]-P_{X}[P[Y\mid A=1,X]]$
$=\psi(P_{X}\times\widehat{P}_{Y,A|X})-\psi(P_{X}\times P_{Y,A|X}).$
For the statistical functional $Q\mapsto\psi(Q)$, let $\varphi$ be its pathwise derivative with respect to $Q_{Y,A|X}$; we normalize it w.l.o.g. so that $Q[\varphi(Z;Q)|X]=0$ for all $Q$. For the ATE (2), it is well-known that
$\varphi(Z;P)=\frac{A}{P[A=1|X]}(Y-P[Y|A=1,X])=\frac{A}{\pi(X)}(Y-\mu(X)).$ (3)
See Kennedy [24] for an accessible primer on such derivations. Then, a (distributional) first-order Taylor expansion gives
[…]
$=-P[\varphi(Z;\widehat{P})]+R_{2}(\widehat{P},P)$ (4)
where $R_{2}$ is a second-order remainder term. Conclude that the first-order error term of the plug-in approach $P_{X}[\widehat{\mu}(X)]$ is given by $-P[\varphi(Z;\widehat{P})]$.
A common approach in semiparametric statistics to better estimate $\psi(P)$ is to explicitly correct for this first-order error term. As we discuss later in Section 6, the first-order correction leads to asymptotic optimality properties like semiparametric efficiency—providing the shortest possible confidence interval—and double robustness—achieving estimator consistency even if only one of $\widehat{\mu}(X),\widehat{\pi}(X)$ is consistent.
One-Step Correction (Augmented Inverse Propensity Weighting) [33, 31]
The most popular first-order correction approach adjusts the plug-in estimator to account for the error term by moving the first-order error term to the left-hand side in the Taylor expansion (4), to achieve second-order error rates
$P\left[\widehat{\mu}(X)+\varphi(Z;\widehat{P})\right]-P[\mu(X)]=R_{2}(\widehat{P},P).$
Using the empirical distribution to approximate the debiased estimator, we arrive at the augmented inverse propensity weighted estimator
$\widehat{\psi}_{\textrm{debiased}}$ $:=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\mu}(X_{i})+\varphi(Z_{i};\widehat{P})\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\mu}(X_{i})+\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\widehat{\mu}(X_{i}))\right).$ (5)
Targeting (Targeted Maximum Likelihood Estimation) [40]
Targeting takes an alternative approach to first-order correction. It commits to the use of the plug-in estimator, and constructs a tailored adjustment to the existing nuisance parameter estimate $\widehat{\mu}(X)$ to set the first-order error to zero. The fluctuation takes place in the direction of a task-specific random variable (“clever covariate”), which takes different forms depending on the estimand [40].
In the case of the ATE with $Y(0):=0$ and general outcomes $Y$ (including continuous and unbounded $Y$), the functional form of the pathwise derivative (3) motivates the use of $\frac{A}{\widehat{\pi}(X)}$ as the clever covariate: targeting uses an adjusted nuisance estimate $\widehat{\mu}(A,X)+\epsilon^{\star}\frac{A}{\widehat{\pi}(X)}$ in place of $\widehat{\mu}(A,X)$ in the plug-in for $\psi(P)=P[\mu(1,X)]$, where $\epsilon^{\star}$ is chosen to solve the targeted maximum likelihood problem
$\epsilon^{\star}:=\mathop{\rm argmin}_{\epsilon\in\mathbb{R}}~{}~{}\frac{1}{n}\sum_{i=1}^{n}A\left(Y-\widehat{\mu}(A,X)-\epsilon\frac{A}{\widehat{\pi}(X)}\right)^{2}.$ (6)
From the KKT conditions, it is easy to see that the solution to this problem removes the finite-sample estimate of the first-order error term (4) of the plug-in estimator relying on $\widehat{\mu}(A,X)+\epsilon^{\star}\frac{A}{\widehat{\pi}(X)}$.
Taking the empirical approximation to the targeted maximum likelihood problem (6), the optimal solution is given by
$\epsilon^{\star}:=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})^{2}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\widehat{\mu}(X_{i})).$ (7)
Thus, we arrive at an explicit formula for the targeted maximum likelihood estimator
$\widehat{\psi}_{\rm targeted}$ $:=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\mu}(X_{i})+\epsilon^{\star}\frac{1}{\widehat{\pi}(X_{i})}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}(X_{i})+\frac{\sum_{i=1}^{n}\frac{1}{\widehat{\pi}(X_{i})}}{\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})^{2}}}\cdot\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\widehat{\mu}(X_{i})).$ (8)
Targeted Regularization [12, 38]
Instead of modifying the outcome model through a post-processing step (6), we can instead regularize the usual training objective for the outcome model using a similar adjustment term, an approach referred to as targeted regularization in the literature [12, 38], where it is demonstrated on e.g. neural networks learned via gradient descent. An additional difference between TMLE and targeted regularization is the data splits involved. In TMLE, the targeting objective (6) is applied to the dataset used for plug-in estimation ($Z_{1:n}$ in the expressions above), while in targeted regularization, the targeting objective is applied to the dataset used for training $\widehat{\mu}$, which has not explicitly described above, as so far, $\widehat{\mu},\widehat{\pi}$ are taken as given. These data splits are often different; see Section 4 for more discussion on data splits, and Section 6 for a recommended data splitting setup.
3 Constrained Learning Framework
The aforementioned approaches to first-order correction take the fitted nuisance estimate as given, and make adjustments to either the estimator (one-step estimation (5)) or the nuisance estimate (targeting (6)). In this section, we propose the constrained learning framework to first-order correction where we train the nuisance parameter to be the best nuisance estimator subject to the constraint that the first-order error term (4) is zero.
Our method, which we call the C-Learner, is a general method for adapting machine learning models to explicitly consider the semiparametric nature of the downstream task during training. For one-step estimation and targeting, in the sections before, we would use a fitted outcome model $\widehat{\mu}$ that minimizes the squared prediction loss, with
$\widehat{\mu}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}P_{\rm train}[A(Y-\tilde{\mu}(X))^{2}],$ (9)
where the loss minimization is performed over the (often separate, auxiliary) training data split $P_{\rm train}$ (which we discuss in more detail in Section 4). In contrast, for C-Learner, we instead solve the constrained optimization problem
$\widehat{\mu}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{\rm train}[A(Y-\tilde{\mu}(X))^{2}]:\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\tilde{\mu}(X_{i}))=0\right\},$ (10)
and our final estimator is a plug-in using nuisance estimator $\widehat{\mu}^{C}$:
$\widehat{\psi}_{\rm C-Learner}:=\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}^{C}(X_{i})$ (11)
so that the constraint in Equation (10) is applied to the same data used in the plug-in (11).
The constraint (10) ensures the corresponding plug-in estimator has a finite-sample estimate of the first-order error term (4) of 0. The training objective for $\widehat{\mu}^{C}$ thus optimizes for the best outcome model fit using the training data, subject to the constraint that the plug-in estimator is asymptotically optimal. In practice, there are several computational approaches to (approximately) solve the stochastic optimization problem (10) depending on the function class $\mathcal{F}$. In Section 4, we describe how to instantiate C-Learner in several different ways: with $\hat{\mu}^{C}$ being linear models, gradient boosted regression trees (XGBoost), and neural networks. We empircally demonstrate these in Section 5.
Like other first-order correction methods, the C-Learner can be implemented with cross-fitting [10], which we discuss further in Sections 4 and 6. Although we illustrate the C-Learner in the ATE setting to simplify the exposition, our estimation approach generalizes to other estimands that are continuous linear functions of the outcome. See Appendix A for a discussion.
3.1 Comparison with Other Approaches to Asymptotic Optimality
The constrained learning framework provides a unifying perspective to existing approaches to first-order correction. First, a variant of AIPW (5), a standard debiased estimator, is a C-Learner over a very restricted class $\mathcal{F}$ of outcome models. For a pre-trained outcome model $\widehat{\mu}$, consider the constrained optimization problem (10) over the model class :=$\mathcal{F}:=\{\widehat{\mu}(X)+\epsilon:\epsilon\in\mathbb{R}\}$. In order to satisfy the constraint (10), we must have $\epsilon^{*}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\left[\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\widehat{\mu}(X_{i}))\right]$, which gives
$\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}^{C}(X_{i})=\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}(X_{i})+\left(\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\left[\frac{A_{i}}{\widehat{\pi}(X_{i})}(Y_{i}-\widehat{\mu}(X_{i}))\right].$ (12)
Compared to the AIPW (5), the above estimator uses a different normalization for the correction term. This normalization term $\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}$ has expectation 1 if we use the true propensity score $\pi$ in place of $\widehat{\pi}$. We refer to this estimator as the self-normalized AIPW.
For general outcome variables (including continuous and unbounded random variables), we can also show that targeting (8) can be viewed as a C-Learner over a limited model class. By inspection, the first-order condition to the TMLE problem (6) is given by
$\frac{1}{n}\sum_{i=1}^{n}\left[\frac{A_{i}}{\widehat{\pi}(X_{i})}\left(Y_{i}-\widehat{\mu}(X_{i})-\epsilon\frac{A_{i}}{\widehat{\pi}(X_{i})}\right)\right]=0.$ (13)
Thus, targeting is a C-Learner where a pre-trained outcome model $\widehat{\mu}(X)$ is perturbed along a specific direction to become $\widehat{\mu}^{C}(X):=\widehat{\mu}(X)+\epsilon\frac{A}{\widehat{\pi}(X)}$. Although contrived, reframing the constrained optimization problem (10) with :=$\mathcal{F}:=\{\widehat{\mu}(X)+\epsilon\frac{A}{\widehat{\pi}(X)}:\epsilon\in\mathbb{R}\}$ thus provides a new way to view TMLE (8).
These existing approaches to first-order correction rely on adding estimated ratios $(A/\widehat{\pi}(X))$ (one-step estimation (5)) or fluctuating the outcome model $\widehat{\mu}$ in a specific direction—along $1/\widehat{\pi}(X)$—in order to achieve unbiased plug-in estimators (targeting (6)). In contrast, our proposed method achieves asymptotic optimality without such limited adjustments to the model class. It simply trains the nuisance parameter $\widehat{\mu}^{C}(\cdot)$ so the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}^{C}(X_{i})$ satisfies the criterion for asymptotic optimality in the most direct way possible, within the originally chosen model class of interest.
In practice, we observe that compared to one-step estimation and targeting, our formulation avoids the potential numerical instability that comes from $1/\widehat{\pi}(X)$ taking extreme values when the estimated treatment probability $\widehat{\pi}(X)$ is small. Plug-in estimators of learned outcome models have been observed to be numerically stable (e.g., [23]) and empirically, the C-Learner inherits these benefits as we showcase Section 5.
4 Methodology
The constrained learning framework can be instantiated in many ways depending on the function class $\mathcal{F}$ for outcome models. Concretely, we illustrate the versatility of the C-Learner by presenting approximate solution methods to the constrained optimization problem (10) for linear models, gradient boosted regression trees, and deep neural networks. We empirically demonstrate these instantiations later in Section 5.
We rely on sample splitting and cross-fitting. Let $P_{\rm train}$ be the split on which nuisance parameters such as outcome model or the propensity score $\widehat{\pi}(\cdot)$ are trained, and let $P_{\rm val}$ be the split on which we perform model selection on them. We use a separate split $P_{\rm eval}$ to evaluate our final causal estimators, as well as assessing their first-order errors. Our constrained learning framework (10) can thus be rewritten as
$\widehat{\psi}_{\rm C-Learner}:=P_{\rm eval}\left[\widehat{\mu}^{C}(X)\right]~{}~{}~{}~{}~{}\mbox{where}$
$\widehat{\mu}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{\rm train}[A(Y-\tilde{\mu}(X))^{2}]:P_{\rm eval}\left[\frac{A}{\widehat{\pi}(X)}(Y-\tilde{\mu}(X))\right]=0\right\},$ (14)
We let $P_{\rm eval}=P_{\rm val}$ except when replicating results from other works.
4.1 Linear Models
When outcome models are linear functions of $X$, the constrained learning problem (14) has an analytic solution. Using $\vec{\cdot}$ to denote stacked observations, define
$\begin{bmatrix}\vec{Y}_{\rm train}\\ \vec{X}_{\rm train}\\ \vec{H}_{\rm train}\end{bmatrix}:=\begin{bmatrix}\{Y_{i}\}_{i\in\mathcal{I}_{\rm train}}\\ \{X_{i}\}_{i\in\mathcal{I}_{\rm train}}\\ \left\{\frac{A_{i}}{\widehat{\pi}(X_{i})}\right\}_{i\in\mathcal{I}_{\rm train}}\end{bmatrix}\qquad\mbox{and}\qquad\begin{bmatrix}\vec{Y}_{\rm eval}\\ \vec{X}_{\rm eval}\\ \vec{H}_{\rm eval}\end{bmatrix}:=\begin{bmatrix}\{Y_{i}\}_{i\in\mathcal{I}_{\rm eval}}\\ \{X_{i}\}_{i\in\mathcal{I}_{\rm eval}}\\ \left\{\frac{A_{i}}{\widehat{\pi}(X_{i})}\right\}_{i\in\mathcal{I}_{\rm eval}}\end{bmatrix}$
where $\mathcal{I}_{\rm train}=\{i\in\mbox{train}:A_{i}=1\}$ and $\mathcal{I}_{\rm eval}=\{i\in\mbox{eval}:A_{i}=1\}$ are indices with observations in each data split. Using this notation, the constrained learning problem (14) can be rewritten as
$\widehat{\theta}^{C}=\mathop{\rm argmin}_{\theta}\left\{\frac{1}{2}\|\vec{Y}_{\rm train}-\vec{X}_{\rm train}\theta\|^{2}:\;\vec{H}_{\rm eval}^{\top}(\vec{Y}_{\rm eval}-\vec{X}_{\rm eval}\theta)\right\}.$
The KKT conditions characterize the primal-dual optimum $(\widehat{\theta}^{C},\widehat{\lambda})$
$\widehat{\theta}^{C}=(\vec{X}_{\rm train}^{\top}\vec{X}_{\rm train})^{-1}\vec{X}_{\rm train}^{\top}(\vec{Y}_{\rm train}+\widehat{\lambda}\vec{H}_{\rm train})~{}~{}\mbox{where}~{}~{}\widehat{\lambda}=\frac{\vec{H}_{\rm eval}^{\top}(\vec{Y}_{\rm eval}-\vec{Y}_{\rm ols})}{\vec{H}_{\rm eval}^{\top}\vec{X}_{\rm eval}(\vec{X}_{\rm train}^{\top}\vec{X}_{\rm train})^{-1}\vec{X}_{\rm train}^{\top}\vec{H}_{\rm train}}$
and $\vec{Y}_{\rm ols}:=\vec{X}_{\rm train}^{\top}(\vec{X}_{\rm train}^{\top}\vec{X}_{\rm train})^{-1}\vec{X}_{\rm train}^{\top}\vec{Y}_{\rm train}$. Note that $\widehat{\theta}^{C}$ is the OLS with respect to the pseudo-label $\vec{Y}_{\rm train}+\widehat{\lambda}\vec{H}_{\rm train}$ and the dual variable shifts the observed outcomes in the direction of $\vec{H}_{\rm train}$ similarly to targeting, but with additional reweighting using covariates.
4.2 Gradient boosted regression trees
We consider outcome models that are gradient boosted regression trees. The gradient boosting framework [17] iteratively estimates the functional gradient $g_{j}$ of the loss function evaluated on the current function estimate $\widehat{\mu}_{j}$. For the standard MSE loss $\ell(\mu;X,A,Y):=A(Y-\mu(X))^{2}$, we use weak learners in $\mathcal{G}$ (e.g., shallow decision trees) to compute
$\widehat{g}_{j+1}\in\mathop{\rm argmin}_{g\in\mathcal{G}}P_{\rm train}[(A(g_{j}(X,Y)-g(X))^{2}]~{}~{}~{}\mbox{where}~{}~{}~{}g_{j}(X,Y;\widehat{\mu}_{j}):=\frac{\partial}{\partial\mu}\ell(\mu;X,Y)\Big{|}_{\mu=\widehat{\mu}_{j}}.$
and set $\widehat{\mu}_{j+1}=\widehat{\mu}_{j}-\eta\widehat{g}_{j+1}$ for some step size $\eta$. As we assume $Y:=0$ in our setting, we let $\mu(X)$, $g_{j}(X,Y)$, and $g(X,Y)$ be shorthand for $\mu(A,X)$, $g_{j}(X,A,Y)$, and $g(X,A,Y)$ when setting $A=1$, and we let $\mu(0,X),g_{j}(0,X),g(0,X)=0$. The process repeats $J$ times until a maximum number of weak learners are fitted or an early stopping criterion is met.
Algorithm 1 C-Learner with Gradient Boosted Regression Trees
1:Input: learning rate $\eta$, max trees $J$ and $K$, $\widehat{\mu}_{0}:=0$
2:for $j=0,2,\dots,J-1$ do
3:     Modify functional gradient $g_{j}=Y-\widehat{\mu}_{j}(X)$ to $\tilde{g}_{j}=g_{j}+\epsilon^{\star}_{j}\cdot A/\widehat{\pi}(X)$ where $\epsilon_{j}^{\star}:=\tfrac{P_{\rm train}[(Y-\widehat{\mu}_{j}(X))\cdot A/\widehat{\pi}(X)]}{P_{\rm train}[A/\widehat{\pi}(X)^{2}]}$ as in (15)
4:     Compute $\widehat{g}_{j}=\mathop{\rm argmin}_{g\in\mathcal{G}}P_{\rm train}[A(\tilde{g}_{j}-g(X))^{2}]$ and update $\widehat{\mu}_{j+1}=\widehat{\mu}_{j}-\eta\widehat{g}_{j}$
5:end for
6:for $k=0,2,\dots,K-1$ do
7:     Compute gradient $\tilde{g}_{k}=\epsilon^{\star}_{J+k}\cdot A/\widehat{\pi}(X)$ where $\epsilon_{J+k}^{\star}=\tfrac{P_{\rm eval}[(Y-\widehat{\mu}_{J+k}(X))\cdot A/\widehat{\pi}(X)]}{P_{\rm eval}[A/\widehat{\pi}(X)^{2}]}$
8:     Compute $\widehat{g}_{k}=\mathop{\rm argmin}_{g\in\mathcal{G}}P_{\rm eval}[A(\tilde{g}_{k}-g(X))^{2}]$ and update $\widehat{\mu}_{J+k+1}=\widehat{\mu}_{J+k}-\eta\widehat{g}_{k}$
9:end for
10:Return final outcome model $\widehat{\mu}^{C}_{\rm XGB}:=\widehat{\mu}_{J+K}$
Constrained gradient boosting
We are interested in minimizing the squared loss subject to the constraint (10). We consider a two stage procedure. First, we perform gradient boosting where instead of the functional gradient of the loss $g_{j}$, we use a modified version
$\tilde{g}_{j}:=g_{j}+\epsilon^{\star}_{j}\cdot\frac{A}{\widehat{\pi}(X)}~{}~{}~{}\mbox{where}~{}~{}~{}\epsilon_{j}^{\star}=\mathop{\rm argmin}_{\epsilon}\left\{P_{\rm train}\left[A\left(Y-\widehat{\mu}_{j}(X)-\epsilon\cdot\frac{A}{\widehat{\pi}(X)}\right)^{2}\right]\right\}$ (15)
given by the targeting objective (6) applied to $\widehat{\mu}_{j}$ on the dataset $P_{\rm train}$. The modification $\tilde{g}_{j}$ allows subsequent weak learners to be fit in a direction that reduces the loss and makes the plug-in estimator closer to satisfying our constraint on $P_{\rm train}$. To ensure that the constraint (14) is satisfied on $P_{\rm eval}$, the second stage fits weak learners to the following gradient of constraint violation
$\tilde{g}_{k}:=\epsilon^{\star}_{J+k}\cdot\frac{A}{\widehat{\pi}(X)}~{}~{}\mbox{where}~{}~{}\epsilon^{\star}_{J+k}=\mathop{\rm argmin}_{\epsilon}\left\{P_{\rm eval}\left[A\left(Y-\widehat{\mu}_{J+k}(X)-\epsilon\cdot\frac{A}{\widehat{\pi}(X)}\right)^{2}\right]\right\}.$
We summarize the method above in pseudo-code in Algorithm 1, which we implement using the XGBoost package with custom objectives  [9]. Hyperparameters for the first stage (learning rate $\eta$, and other properties of the weak learners such as max tree depth) are selected to have the lowest MSE loss on $P_{\rm val}$. Other hyperparameters such as max number of trees $J$ and $K$, may be set on $P_{\rm val}$, or alternatively, by early stopping, with evaluation on different splits within $P_{\rm train}$. These hyperparameters are re-used in the second stage.
4.3 Neural Networks
When outcome models are neural networks $\widehat{\mu}_{\theta}(x)$ with weights $\theta$, we consider the usual MSE loss with a Lagrangian regularizer for the constraint (14)
$R(\theta)=P_{\rm train}\left[A(Y-\widehat{\mu}_{\theta}(X))^{2}\right]+\lambda\cdot P_{\rm eval}\left[\frac{A}{\widehat{\pi}(X)}(Y-\widehat{\mu}_{\theta}(X))\right]^{2}.$
We optimize the above objective using stochastic gradient descent, where we take mini-batches of the training set to approximate the gradient of the first term and take a full-batch gradient on the evaluation set for the second term. At the end of every training epoch, the constraint on $P_{\rm eval}$ is enforced exactly by adjusting the constant bias term $\theta_{\rm bias}$ in the neural network:
$\theta_{\rm bias}\leftarrow\theta_{\rm bias}+\left(P_{\rm eval}\left[\frac{A}{\widehat{\pi}(X)}\right]\right)^{-1}P_{\rm eval}\left[\frac{A}{\widehat{\pi}(X)}(Y-\widehat{\mu}_{\theta}(X))\right].$ (16)
We choose the early stopping epoch to be the one that minimizes MSE loss on $P_{\rm val}\left[A(Y-\widehat{\mu}_{\theta}(X))^{2}\right]$. For choosing hyperparameters (e.g. learning rate, $\lambda$), we consider two options: (1) choosing the hyperparameters that minimize MSE loss on $P_{\rm val}$, and (2) choosing hyperparameters (among ones with reasonable MSE loss on $P_{\rm val}$ that minimize the size of the bias shift (16) in the first epoch, as a small bias shift indicates a regularizer that is successful. While the first method is more standard, the second more closely encourages satisfying the constraint over the course of training, to avoid big jumps in optimization. We note that model selection for nuisance parameters is an active area of research [18, 35, 37]. We explore these in Section 5.2.
5 Experiments
In this section, we present a series of experiments to demonstrate the flexibility and empirical performance of the C-Learner. In Section 5.1, we consider well-studied tabular simulated settings where one-step estimation and targeting methods perform poorly [23, 30], and we observe that the C-Learner, implemented with linear models and gradient boosted regression trees, performs better in comparison. To test the scalability of our approach, in Section 5.2 we construct and study a high-dimensional causal inference setting with text features. We fine-tune a language model [36] under our constrained learning framework. In both settings, we observe that C-Learner outperforms one-step estimation and targeting. Lastly, in Appendix C.3, we study a common tabular dataset (Infant Health and Development Program [6]); the C-Learner implemented with gradient boosted regression trees matches the performance of one-step estimation and targeting.
5.1 Kang & Schafer Dataset [23] (tabular)
We start with the synthetic tabular setting constructed by Kang and Schafer [23], who demonstrate empirically that the direct method (“OLS” plug-in estimator with a linear outcome model) achieves better performance than asymptotically optimal methods in their setting. In response, Robins et al. [30] note that settings in which some subpopulations have a much higher probability of being treated than others [34, 32, 23] can be very challenging for asymptotically optimal methods, as such methods are sensitive to propensity weights that can take extreme values in these settings. Thus, investigating the performance of the C-Learner in this challenging setting contributes to understanding its empirical stability on a well-known dataset.
Our goal is to estimate $\psi:=P[Y(1)]=P[P[Y|A=1]]=P[\mu(X)]$ from data $(X,A,AY)$ (assuming $Y(0):=0$). The true outcome and treatment mechanisms to depend on covariates $\xi\sim N(0,I)\in\mathbb{R}^{4}$ and $\varepsilon\sim N(0,1)$
$Y$ $=210+27.4\xi_{1}+13.7\xi_{2}+13.7\xi_{3}+13.7\xi_{4}+\varepsilon,$
$\pi(\xi)$ $=\frac{\exp(-\xi_{1}+0.5\xi_{2}-0.25\xi_{3}-0.1\xi_{4})}{1+\exp(-\xi_{1}+0.5\xi_{2}-0.25\xi_{3}-0.1\xi_{4})}.$
Here, $P[Y(1)]=P[Y\mid A=1]=200$ and $P[Y]=210$ so that a naive average of the treated units is biased by -10. For a random sample of 100 data points, the true propensity score can be as low as 1% and as high as 95%. Instead of observing $\xi$, the modeler observes
$X_{1}=\exp(\xi_{1}/2),\;X_{2}=\xi_{2}/(1+\exp(\xi_{1}))+10,\;X_{3}=(\xi_{1}\xi_{3}/25+0.6)^{3},\;X_{4}=(\xi_{2}+\xi_{4}+20)^{2}.$
Next, we demonstrate the instantiations of the C-Learner with two different function classes: linear function classes and gradient boosted regression trees.
Linear Outcome Models.
Following Kang and Schafer [23], we consider outcome models that are linear in $X$ and propensity models that are linear logistic regressions of the treatment on $X$. Therefore, both models are misspecified, as is common in real world applications. We use squared loss to train the outcome model on treated units ($A=1$) and we use the logistic loss to learn the propensity (treatment) model. This leads to outcome models with $R^{2}$ close to 0.99 and propensity models with ROC around 0.75. Following Kang & Schafer, we do not distinguish between $P_{\rm train}=P_{\rm val}=P_{\rm eval}$.
We present a comparison of methods in Table 1. In this table, “Direct” refers to the plug-in with the outcome model trained as usual (“OLS” in [23]), “IPW-POP” [23] refers to self-normalized IPW and “AIPW-POP” refers to self-normalized AIPW. Observe that the C-Learner outperforms other asymptotically optimal methods in terms of mean absolute error (MAE). [^1] The median absolute error is reported in [23]. Qualitative results are the same under this additional metric, which we omit for easier exposition. In Appendix C.1.2, we also show that this is the case for other metrics such as RMSE and Median Absolute Error. Our method is the only asymptotically optimal method that is comparable to the direct method in this setting, indicating strong numerical stability, despite extreme inverse propensity weights. In some cases, the C-Learner improves upon AIPW and TMLE by orders of magnitude.
(a) $N=200$
Method Bias Mean Abs Err
Direct -0.00 (0.10) 2.60 (0.06)
IPW 22.10 (2.58) 27.28 (2.53)
IPW-POP 3.36 (0.29) 5.42 (0.26)
AIPW -5.08 (0.474) 6.16 (0.46)
AIPW-POP -3.65 (0.20) 4.73 (0.18)
TMLE -111.59 (41.07) 112.15 (41.07)
C-Learner -2.45 0.12) 3.57 (0.09)
(b) $N=1000$
Bias Mean Abs Err
-0.43 (0.04) 1.17 (0.03)
105.46 (59.84) 105.67 (59.84)
6.83 (0.33) 7.02 (0.32)
-41.37 (24.82) 41.39 (24.82)
-8.35 (0.43) 8.37 (0.43)
-17.51 (3.49) 17.51 (3.49)
-4.40 (0.077) 4.42 (0.07)
Table 1: Comparison of estimator performance on misspecified datasets from Kang and Schafer [23] in 1000 tabular simulations. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing asymptotically optimal method in bold. Standard errors are displayed within parentheses to the right of the point estimate.
As methods such as IPW and AIPW are known to be unstable if the propensity score $\widehat{\pi}(X)$ is close to 0, a common ad-hoc heuristic is to clip $\widehat{\pi}(X)$ for a chosen small $\eta>0$ so that if $\widehat{\pi}(X)<\eta$, we redefine $\widehat{\pi}(X):=\eta$. Dealing with extreme $\widehat{\pi}(X)$ is an an active area of research; choosing such an $\eta$ is an art, and some works propose to avoid these difficult $X$ entirely by adjusting the estimand to exclude or down-weight those low-overlap covariates [14, 25]. We show in Appendix C.1.1 that, notably, clipping seems essential for asymptotically optimal methods AIPW and TMLE to perform well.
While the direct method performs extremely well on this dataset, this estimator performs poorly on the flipped version of this task in which we would like to estimate $\psi^{\prime}:=P[Y(0)]$ instead. In their comment to Kang and Schafer [23], Robins et al. [30] thus asks the following question (rewritten to match our notation): “Can we find doubly-robust estimators that, under the authors’ chosen joint distribution for $(X,A,(1-A)Y)$, both perform almost as well as the direct method for $P[Y(1)]$ and yet perform better than the direct method for $P[Y(0)]$?” We observe that the C-Learner is such an estimator. In Appendix C.1.2, we show that, indeed, when estimating $P[Y(0)]$, the performance of all the asymptotically efficient methods are equivalent and superior to the direct method, and in the original version for estimating $P[Y(0)]$, C-Learner had the best performance across all estimators other than the direct method.
Gradient Boosted Regression Tree Outcome Models.
We now demonstrate the flexibility of the C-Learner by using gradient boosted regression trees as outlined in Section 4.2. As before, the propensity model $\widehat{\pi}$ is fit as a logistic regression on covariates $X$. Since sample splitting is more appropriate for this flexible model class, we use with cross-fitting as we describe in Section 6, with $K=2$ folds. Additional implementation details, such as the grid of hyperparameters for tuning and coverage results are deferred to Appendix C.1.1.
The results for the mean absolute error and their respective standard errors are displayed in Table 2. The C-Learner outperforms the direct method and achieves the best mean absolute error (MAE) among all estimators and across all sample sizes. For comparison, we also include a plug-in method where the outcome model is learned using only the first stage in Section 4.2. In the results in Table 2 it is labeled as “Lagrangian”, as it can be seen as a Lagrangian relaxation of the C-Learner framework. Especially for small sample sizes (e.g. $N=200$), the IPW, AIPW, and TMLE estimators perform very poorly, likely due to being sensitive to extreme propensity weights. Self-normalization for both IPW (“IPW-POP”) and AIPW (“AIPW-POP”) is crucial in these settings.
In Appendix C.1.1, we also present the results when arbitrary values of clipping of 0.1% and 5% are used. In this case, C-Learner again performs better (clipping 0.1%) or equivalently to TMLE and AIPW-POP (clipping 5%) and is superior to other methods.
(a) $N=200$
Method Bias Mean Abs Err
Direct -5.12 (0.10) 5.30 (0.09)
IPW 1192 (919) 1206 (919)
IPW-POP -1.01 (0.10) 7.29 (0.33)
Lagrangian -4.44 (0.10) 3.57 (0.09)
AIPW 275 (215) 280 (215)
AIPW-POP -0.82 (0.24) 4.53 (0.19)
TMLE 487 (345) 10927 (493)
C-Learner -2.89 (0.10) 3.53 (0.07)
(b) $N=1000$
Bias Mean Abs Err
-3.48 (0.04) 3.49 (0.04)
28.3 (3.23) 32.30 (3.23)
2.26 (0.29) 4.85 (0.26)
-2.29 (0.04) 2.34 (0.05)
2.28 (0.52) 4.58 (0.51)
0.36 (0.14) 2.74 (0.15)
17.3 (10.20) 20.05 (10.23)
-1.92 (0.04) 2.03 (0.04)
Table 2: Comparison of estimator performance on misspecified datasets from Kang and Schafer [23] in 1000 tabular simulations using gradient boosted regression trees. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing asymptotically optimal method in bold. Standard errors are displayed within parentheses to the right of the point estimate.
5.2 CivilComments Dataset (NLP)
Neural networks can learn good feature representations for image and text data. Despite the recent interest in using neural networks for fitting nuisance parameters [12, 38], prior work do not evaluate their methods outside of tabular settings. To fill this gap, we construct a semisynthetic causal inference dataset using text covariates.
Setting
Content moderation is a fundamental problem for maintaining the integrity of social media platforms. We consider a setting in which we wish to measure the average level of toxicity across all user comments. It is infeasible to have human experts label all comments for toxicity, and the comments that get flagged for human labeling may be a biased sample. We pose this as a missing data problem, which is equivalent to our usual goal of estimating the ATE (2) as in Section 2.1.
We use the CivilComments dataset [13], which contains real-world online comments and corresponding human-labeled toxicity scores. The dataset contains toxicity labels $Y(1)$ for all comments $X$ (which provides a ground truth to compare to), and we construct the labeling (treatment) mechanism $A\in\{0,1\}$ to induce selection bias. Specifically, whether the toxicity label for a comment can be observed is drawn according to the Bernoulli random variable
$A\sim\textrm{Bernoulli}(g(X))~{}~{}~{}\mbox{where}~{}~{}~{}g(X)=\textrm{clip}(b(X),l,u),\quad\textrm{clip}(y,l,u)=\begin{cases}l&\textrm{if}\;y<l\\ y&\textrm{if}\;l\leq y\leq u\\ u&\textrm{if}\;y>u\end{cases}.$
In practice, we set $u=0.9$ and $l=10^{-4}$. Note that a lower $l$ implies less overlap and more extreme propensity weights. Here, $b(X)\in[0,1]$ is a continuous measure of whether comment $X$ relates to the demographic identity “Black”, as provided available in the dataset. The labeled data suffers from selection bias: comments mentioning the demographic identity “Black” are positively correlated to toxicity within this dataset. Therefore, a naive average of toxicity over labeled units would overestimate the overall toxicity. From a causal perspective, we have confounding, induced by selection bias, which we can handle by using estimators for the ATE.
Procedure
We demonstrate how C-Learner can be instantiated with neural networks. To learn $\widehat{\pi}$, $\widehat{\mu}$, and $\widehat{\mu}_{C}$, we fine-tune a pre-trained DistilBERT model [36] with a linear head using stochastic gradient descent, with $\widehat{\mu}^{C}$ learned using the procedure described in Section 4.3. We train $\widehat{\mu}$ using squared loss on labeled units, and propensity models $\widehat{\pi}$ using the logistic loss.
We construct nuisance estimators on 100 re-drawn datasets of size 2000 each, from the full dataset of size 405,130. On each dataset draw, we use cross-fitting, as described in Section 6, with $K=2$ folds, for all estimators. We consider two ways of picking hyperparameters ($\lambda$ and learning rate), as described in Section 4.3: one in which we choose hyperparameters for the best mean squared error on validation data, and one in which we choose hyperparameters that minimize the size of the bias shift. Additional experiment details are in Appendix C.2.1.
Results
Method Bias Mean Abs Err
Direct 0.173 (0.008) 0.177 (0.007)
IPW 0.504 (0.084) 0.546 (0.081)
IPW-POP 0.114 (0.017) 0.153 (0.014)
AIPW 0.084 (0.043) 0.307 (0.032)
AIPW-POP 0.116 (0.018) 0.161 (0.014)
TMLE -1.264 (1.361) 1.802 (1.355)
C-Learner (best val MSE) 0.103 (0.015) 0.141 (0.011)
C-Learner (smallest bias shift) 0.075 (0.012) 0.115 (0.008)
Table 3: Comparison of estimators in the CivilComments [13] semi-synthetic dataset over 100 re-drawn datasets, with $l=10^{-4}$. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing method in bold. Standard errors are displayed within parentheses to the right of the point estimate.
In Table 3, we find that C-Learner with both variants for hyperparameter section has the best mean absolute error compared to other methods. In this setting, C-Learner beats both not only other asymptotically optimal methods, but all methods. We observe that C-Learner performs better compared to other asymptotically optimal methods in datasets with low overlap in the data generating process (lower $l$; we investigate $l=10^{-2},10^{-3},10^{-4}$) in Appendix C.2.2.
The results in this NLP setting echo the results in our tabular experiments, suggesting that C-Learner’s reliable performance extends to settings with more complex covariates and where nuisance parameters are modeled as neural networks trained with standard stochastic optimization methods.
6 Asymptotic Optimality
In this section, we identify conditions under which the C-Learner is semiparametrically efficient and doubly robust. In particular, our theoretical treatment provides the asymptotic optimality of one-step estimation (self-normalized AIPW) and targeting (TMLE with unbounded continuous outcomes) since they are also C-Learners satisfying these conditions. We focus on the cross-fitted formulation of the C-Learner. We split the data so that nuisance estimators are fitted on a training (auxiliary) fold and evaluated to form a causal estimator on an evaluation (main) fold. To utilize the entire dataset, we evaluate over the entire dataset, over multiple splits [42, 10].
Formally, we divide a dataset of size $n$ into $K$ disjoint cross-fitting splits. Assuming $K$ evenly divides $n$ for simplicity, let $P_{k,n}$ be the empirical measure over data from the $k$-th split, and $P_{-k,n}$ be the empirical measure over data from all other splits. For each $k=1,\ldots,K$, on the training fold $P_{-k,n}$, we train a propensity score model $\widehat{\pi}_{-k,n}(x)$ to estimate $P[A\mid X=x]$. The C-Learner optimizes the prediction loss evaluated under $P_{-k,n}$, subject to the first-order correction constraint evaluated on the evaluation fold $P_{k,n}$:
$\widehat{\mu}_{-k,n}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{-k,n}[A(Y-\tilde{\mu}(X))^{2}]:P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\tilde{\mu}(X))\right]=0\right\}.$ (17)
The final C-Learner estimator is given by
$\widehat{\psi}^{C}_{n}:=\frac{1}{K}\sum_{k=1}^{K}P_{k,n}[\widehat{\mu}_{-k,n}^{C}(X)].$ (18)
Using the terminology in Section 4, for fold $k$, we let $P_{\rm train}=P_{-k,n}$, and $P_{\rm val}=P_{\rm eval}=P_{k,n}$.
In addition to its practical benefits, cross-fitting simplifies the proof of asymptotic optimality, especially when nuisance parameter models can be large and complex. (When nuisance model classes are simpler, asymptotic optimality can be shown without explicit sample splitting [24, 39, 43].) In contrast to standard cross-fitting, the constraint (17) is over $P_{k,n}$, which is the same as the data on which it is evaluated in Equation (18). We shortly identify sufficient conditions that guarantee asymptotic optimality (Assumption D).
We require standard conditions on the nuisance parameters $\widehat{\pi}_{-k,n},\widehat{\mu}^{C}_{-k,n}$ in order to show asymptotic optimality.
Assumption A (Overlap).
For some $\eta>0$ and for all $k,n$, we have
[…]
Assumption B (Convergence rates of propensity and constrained outcome models).
For all $k\in[K]$, both $\widehat{\pi},\widehat{\mu}^{C}$ are consistent,
$\|\widehat{\pi}_{-k,n}-\pi\|_{L_{2}(P)}=o_{p}(1),\quad\|\widehat{\mu}_{-k,n}^{C}-\mu\|_{L_{2}(P)}=o_{p}(1)$
and also
$\|\widehat{\pi}_{-k,n}-\pi\|_{L_{2}(P)}\cdot\|\widehat{\mu}_{-k,n}^{C}-\mu\|_{L_{2}(P)}=o_{p}(n^{-\frac{1}{2}}).$
As we discuss below, we can relax these assumptions when guaranteeing double robustness (consistency under misspecified nuisance parameters). As is typical, we assume that outcomes do not differ too much from their means, conditional on covariates.
Assumption C (Outcomes are close to their conditional means).
For all $k,n$, and for some $0<B<\infty$,
$\|A(Y-P[Y\mid X])\|_{L_{2}(P)}=\|A(Y-\mu(X))\|_{L_{2}(P)}\leq B.$
To show the validity of the distributional Taylor expansion (4) for the C-Learner, we require the following condition on the C-Learner outcome model.
Assumption D (Empirical process assumption).
$(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{C}(X)-\mu(X))=o_{p}(n^{-1/2}).$
In Section 3.1, we showed how versions of one-step estimation and targeting can also be considered C-Learners. We show in Appendix B.4 and Appendix B.5 that cross-fitted versions of the self-normalized AIPW and TMLE for continuous and unbounded outcomes, under standard assumptions on $\widehat{\pi}_{-k,n},\widehat{\mu}_{-k,n}$ (with $\widehat{\mu}_{-k,n}$ the usual outcome model learned using Equation (9)) also satisfy D. Thus, the C-Learner results in the Theorems 12 to come apply directly to these versions of one-step estimation and targeting as well.
Under these assumptions, the C-Learner (18) enjoys the following asymptotics; see Appendix B.1 for the proof.
Theorem 1.
Under Assumptions A, B, C, and D,
$\sqrt{n}(\widehat{\psi}_{n}^{C}-\psi(P))\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}N(0,\sigma^{2})~{}~{}~{}\mbox{where}~{}~{}~{}\sigma^{2}:={\rm Var}_{P}\left(\frac{A}{\pi(X)}(Y-\mu(X))+\mu(X)\right).$
Since $\sigma^{2}$ is the semiparametric efficiency bound [21], the C-Learner (18) achieves the tightest confident interval and is optimal in the usual local asymptotic minimax sense [43, Theorem 25.21].
Double robustness
By virtue of their first-order correction, standard approaches like one-step estimation and targeting enjoy double robustness: if either of the propensity model or the outcome model are consistent, then the resulting estimator is consistent. We can show a similar guarantee for the C-Learner. Here, we assume only that either $\widehat{\pi}_{-k,n}$ $\widehat{\mu}_{-k,n}^{C}$ is consistent.
Assumption E (At least one of $\widehat{\pi},\widehat{\mu}^{C}$ is consistent).
For all $k$, the product of the errors for the outcome and propensity models decays as
$\|\widehat{\pi}_{-k,n}-\pi\|_{L_{2}(P)}\cdot\|\widehat{\mu}^{C}_{-k,n}-\mu\|_{L_{2}(P)}=o_{P}(1).$
Using Assumption E in place of Assumption B, we arrive at the following result; see Appendix B.2 for the proof.
Theorem 2.
The C-Learner (18) is consistent under Assumptions A, C, D, E.
7 Discussion
We introduced a constrained learning perspective to causal estimation and semiparametric inference. We pose asymptotically optimal plug-in estimators as those whose nuisance parameters are solutions to a optimization problem, under the constraint that the first-order error of the plug-in estimator with respect to the nuisance parameter estimate is zero. This perspective encompasses versions of one-step estimation and targeting, two existing approaches to achieving asymptotically optimal estimators.
The constrained learning framework enables the development of a new method which we call the Constrained Learner, or C-Learner, which solves this constrained optimization problem in the most direct way. We demonstrate C-Learner’s versatility by instantiating it diverse model classes including linear models, gradient boosted trees, and neural networks. On datasets with tabular and text covariates, we observe the C-Learner performs better than other asymptotically optimal methods in settings with low overlap, and similarly in other settings. Building a theoretical understanding of precisely when and why C-Learner works better than one-step estimation and targeting is a fruitful direction of future work.
We hope this work spurs further investigation of how constrained optimization algorithms [28] can provide more robust causal estimators. Alternative ways to instantiate C-Learner using gradient boosted regression trees include, for example, alternating direction method of multipliers (ADMM) [5]. However, in practice, we found that implementing ADMM in conjunction with XGBoost was enormously slower than Algorithm 1.
Appendix A Extending C-Learner to other estimands
We briefly sketch how the C-Learner can be extended to other estimands, beyond just the ATE with $Y(0)=0$ as in Section 2.1. Let $Z:=(W,Y)\sim P$ and let the target functional $\psi(P)$ be continuous and linear in $\mu(W)=P(Y\mid W)$, the conditional distribution of $Y$ given $W$. For example, we let $W=(X,A)$ in the ATE setting. For other functionals that admit a distributional Taylor expansion with pathwise derivative with respect to $P_{Y,W|X}$, $\varphi(Z)$, then we can similarly formulate the C-Learner again as learning the best $\widehat{\mu}$, subject to the constraint that the estimate of the first-order error term is 0.
When $\psi(P)$ is continuous and linear in $\mu$, by the Riesz representation theorem, if $\psi(P)$ is $L_{2}(P)$-continuous in $\mu$ (see for example Equation (4.4) from [27]), then there exists $a\in L_{2}(P)$ such that for all $\mu\in L_{2}(P)$,
$\psi(P)=P[a(W)\mu(W)].$
This $a(\cdot)$ is commonly referred to as the Riesz representer [11], and the corresponding random variable $a(W)$ can be referred to as the clever covariate [40]. These linear functionals satisfy the following mixed bias property:
$\psi(\widehat{P})-\psi(P)+P[\widehat{a}(W)(Y-\widehat{\mu}(W))]=P[(\widehat{a}(W)-a(W))(\widehat{\mu}(W)-\mu(W))],$
where the first-order term in the distributional Taylor expansion as discussed in Section 2.2 is given by $P[\widehat{a}(W)(Y-\widehat{\mu}(W))]$. We refer the reader to [11] or Proposition 4 of [27] for a discussion about this mixed bias property. Therefore, the C-Learner can be formulated more generally as
$\widehat{\mu}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{\rm train}[\ell(W,Y;\tilde{\mu})]:P_{\rm eval}[\widehat{a}(W)(Y-\tilde{\mu}(W))]=0\right\},$
where $\ell$ is an appropriate loss function for the outcome model.
Below, we provide several specific examples demonstrating how C-Learner could be adapted to various target functionals when $W=(X,A)$.
Average Treatment Effect
We have seen the mean missing outcome setting (10) for estimating the target functional $\psi(P)=P[\mu(X)]$ where $\mu(x):=P[Y\mid A=1,X=x]$. The loss function is $\ell(W,Y;\widehat{\mu})=A(Y-\widehat{\mu}(X))^{2}$. The Riesz representer is $a(X,A)=A/\pi(X)$.
If we no longer assume that $Y(0)=0$ and we are interested in estimating the standard average treatment effect
$\psi(P)=P[Y(1)-Y(0)]=P\left[P[Y\mid A=1,X]\right]-P\left[P[Y\mid A=0,X]\right],$
then the Riesz representer, constrained outcome model, and estimator are, respectively,
$a(W)=\frac{A}{\pi(X)}-\frac{1-A}{1-\pi(X)},$
$\widehat{\mu}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{\rm train}[\ell(X,A,Y;\tilde{\mu})]:P_{\rm eval}\left[\left(\frac{A}{\widehat{\pi}(X)}-\frac{1-A}{1-\widehat{\pi}(X)}\right)(Y-\tilde{\mu}(X,A))\right]=0\right\},$
$\widehat{\psi}^{C}:=P_{\rm eval}\left[\widehat{\mu}^{C}(X,1)-\widehat{\mu}^{C}(X,0)\right].$
Average Policy Effect
For off-policy evaluation, the goal is to optimize over assignment policies $c(X)\in\{0,1\}$ to maximize the expected reward under the policy, using observational data collected under an unknown policy $\pi(x):=P(A=1\mid X)$. Assume the usual causal inference assumptions (SUTVA, ignorability, overlap in Section 2.1). Fixing $c(X)$, the average policy effect is
$\psi(P)$ $=P[c(X)Y(1)+(1-c(X))Y(0)]$
$=P[c(X)P[Y(1)\mid X]+(1-c(X))P[Y(0)\mid X]]$
$=P[c(X)P[Y\mid A=1,X]]+(1-c(X))P[Y\mid A=0,X]]$
Here, the Riesz representer, constrained outcome model, and estimator are, respectively,
$a(X,A)=c(X)\frac{A}{\pi(X)}+(1-c(X))\frac{1-A}{1-\pi(X)},$
$\widehat{\mu}^{C}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{\rm train}[\ell(X,A,Y;\tilde{\mu})]:P_{\rm eval}\left(c(X)\frac{A}{\pi(X)}+(1-c(X))\frac{1-A}{1-\pi(X)}\right)(Y-\tilde{\mu}(X,A))=0\right\},$
$\widehat{\psi}^{C}:=P_{\rm eval}\left[c(X)\widehat{\mu}^{C}(X,1)+(1-c(X))\widehat{\mu}^{C}(X,0)\right].$
Appendix B Proofs
Here, we prove results in Section 6. We show that C-Learner is semiparametrically efficient (Appendix B.1) and doubly robust (Appendix B.2), under assumptions in Section 6. We also show that versions of one-step estimation methods (AIPW) and targeting methods (TMLE) can satisfy these assumptions (Appendix B.4, Appendix B.5), so that semiparametric efficiency and double robustness hold for them immediately as well.
B.1 Proof of Theorem 1
Our proof follows a standard argument. We first use the distributional Taylor expansion (19) to rewrite the estimation error $\widehat{\psi}_{n}^{C}-\psi$ as the sum of three terms. Then, we address these terms one by one, and we will show how only one of these terms contributes to asymptotic variance.
Let $Z=(X,A,Y)$ as defined in Section 2.1. Let $P$ denote the true population distribution of $Z$. Functionals $\psi$ may admit a distributional Taylor expansion, also known as a von Mises expansion [16], where for any distributions $P,\bar{P}$ on $Z$, we can write
[…]
When $\psi$ is the ATE as in our setting, $\psi$ admits such an expansion [21]
$\varphi(Z;\bar{P}):=\frac{A}{\bar{\pi}(X)}(Y-\bar{\mu}(X))+\bar{\mu}(X)-\psi(\bar{P}),$ (20)
where $\bar{\pi}(x):=\bar{P}[A=1\mid X]$ and $\bar{\mu}(x):=\bar{P}[Y\mid A=1,X]$. In particular, we get the following explicit formula for the second-order term
[…]
We will apply this to our C-Learner estimator as defined in Section 6. Recall that $\widehat{\pi}_{-k,n}$ is trained to predict treatment $A$ given $X$ using $P_{-k,n}$, and $\widehat{\mu}_{-k,n}^{C}$ is trained to predict outcome given $X$ and $A=1$ on $P_{-k,n}$, under the constraint that $P_{k,n}[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}^{C}(X)]=0$. Our C-Learner estimator is the mean of plug-in estimators across folds: for each fold, write $\widehat{\psi}^{C}_{k,n}=\psi(\widehat{P}_{k,n}^{C})$ so the C-Learner estimate is the average $\widehat{\psi}^{C}_{n}=\frac{1}{K}\sum_{k=1}^{K}\widehat{\psi}^{C}_{k,n}$.
Noting that any distribution decomposes $\bar{P}=\bar{P}_{X}\times\bar{P}_{A\mid X}\times\bar{P}_{Y\mid A,X}$ and $\psi(\bar{P})=\bar{P}_{X}[\mu(X)]$, the following definitions
$\widehat{P}^{C}_{X;k,n}$ $:=P_{X;k,n}$
$\widehat{P}^{C}_{A\mid X;k,n}[A=1\mid X=x]$ $:=\widehat{\pi}_{-k,n}(x)$
$\widehat{P}^{C}_{Y\mid A,X;k,n}[Y\mid A=1,X=x]$ $:=\widehat{\mu}_{-k,n}^{C}(x)$
provide a well-defined joint distribution $\widehat{P}^{C}_{k,n}$.
For each data fold $k$, we use the distributional Taylor expansion above, where we replace $\bar{P}$ with the joint distribution $\widehat{P}^{C}_{k,n}$
$\psi(\widehat{P}^{C}_{k,n})-\psi(P)$ $=-P\varphi(Z;\widehat{P}^{C}_{k,n})+R_{2}(\widehat{P}^{C}_{k,n},P)$
$=(P_{k,n}-P)\varphi(Z;P)-P_{k,n}\varphi(Z;\widehat{P}^{C}_{k,n})$
$\quad+(P_{k,n}-P)(\varphi(Z;\widehat{P}^{C}_{k,n})-\varphi(Z;P))+R_{2}(\widehat{P}^{C}_{k,n},P).$ (22)
Observe that by using Equation (20) and the definition of the C-Learner,
$P_{k,n}\varphi(Z;\widehat{P}^{C}_{k,n})$ $=P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}^{C}_{-k,n}(X))+\widehat{\mu}^{C}_{-k,n}-\psi(\widehat{P}^{C}_{k,n})\right]$
$=\underbrace{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}^{C}_{-k,n}(X))\right]}_{=0\text{by C-Learner constraint}}+\underbrace{P_{k,n}[\widehat{\mu}^{C}_{-k,n}]-P_{k,n}[\widehat{\mu}^{C}_{-k,n}]}_{=0}=0.$
Taking the average of Equation (B.1) over $k=1,\ldots,K$, we can write the error $\widehat{\psi}^{C}_{n}-\psi$ as the sum of three terms.
$\widehat{\psi}^{C}_{n}-\psi$ $=\frac{1}{K}\sum_{k=1}^{K}\underbrace{(P_{k,n}-P)\varphi(Z;P)}_{S^{*}_{k}}$
$\qquad+\frac{1}{K}\sum_{k=1}^{K}\underbrace{(P_{k,n}-P)\left(\varphi(Z;\widehat{P}^{C}_{k,n})-\varphi(Z;P)\right)}_{T_{1k}}+\frac{1}{K}\sum_{k=1}^{K}\underbrace{R_{2}(\widehat{P}^{C}_{k,n},P)}_{T_{2k}}.$ (23)
Using the decomposition (23), we write
$S^{*}=\frac{1}{K}\sum_{k=1}^{K}S^{*}_{k},\hskip 10.00002ptT_{1}=\frac{1}{K}\sum_{k=1}^{K}T_{1k},\hskip 10.00002ptT_{2}=\frac{1}{K}\sum_{k=1}^{K}T_{2k},$ (24)
so that $\widehat{\psi}^{C}_{n}-\psi=S^{*}+T_{1}+T_{2}$. We address the terms $S^{*},T_{1}$ and $T_{2}$ separately. The first term can be rewritten as
$S^{*}=\frac{1}{K}\sum_{i=1}^{K}(P_{k,n}-P)\varphi(Z;P)=(P_{n}-P)\varphi(Z;P)$
so that by the central limit theorem,
$\sqrt{n}S^{*}\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}{N}\left(0,{\rm Var}(\varphi(Z;P))\right).$
Observe that this quantity depends only on $\psi$ and $P$, so that it cannot be made smaller by choice of estimator. If the variance of an estimator for $\psi$ is ${\rm Var}(\varphi(Z;P))$, then it is semiparametrically efficient in the local asymptotic minimax sense (Theorem 25.21 of [43]). Thus, it suffices to show that the rest of the terms, $T_{1}$ and $T_{2}$, are $o_{P}(n^{-1/2})$, so that
$\sqrt{n}(\widehat{\psi}^{C}-\psi)=\sqrt{n}S^{*}+o_{P}(1)\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}{N}\left(0,{\rm Var}(\varphi(Z;P))\right).$
For a fixed $k$, $|T_{1k}|=o_{P}(n^{-1/2})$ by Assumption D so that $|T_{1}|=o_{P}(n^{-1/2})$ as desired. The second-order remainder term in the distributional Taylor expansion (19) where we replace $\bar{P}$ with $\widehat{P}^{C}_{k,n}$ is
[…]
Under the overlap assumption (Assumption A) and Cauchy-Schwarz,
[…]
$\leq\frac{1}{\eta}\left\|\widehat{\pi}_{-k,n}-\pi\right\|_{L_{2}(P)}\left\|\widehat{\mu}^{C}_{-k,n}-\mu\right\|_{L_{2}(P)}.$ (25)
By Assumption B, $|T_{2k}|=o_{P}(n^{-1/2})$ so that $|T_{2}|=o_{P}(n^{-1/2})$ as desired.
B.2 Proof of Theorem 2
We briefly sketch the proof as it is a minor modification of our previous proof in Section B.1. To show the C-Learner estimator $\widehat{\psi}^{C}_{n}$ is consistent (rather than that our estimator has the desired asymptotics) under Assumption (E) (rather than Assumption (B)), we again rewrite the error $\widehat{\psi}_{n}^{C}-\psi$ as a sum of three terms and show each converges to 0 in probability:
  • $S^{*}$: this term converges to 0 in probability.
  • $T_{2k}$: This also converges to 0 in probability by the same logic as before. Note we only need this to be $o_{P}(1)$ and not $o_{P}(n^{-1/2})$ as we would have required for efficiency.
  • $T_{1k}$: this converges to 0 in probability by Assumption D.
B.3 Showing self-normalized AIPW and TMLE satisfy C-Learner conditions for Theorems (1) and (2)
As self-normalized AIPW and TMLE involve simple adjustments to the unconstrained outcome models, in this section, we state assumptions on the unconstrained outcome model $\widehat{\mu}_{-k,n}$, fitted in the usual manner on the auxiliary fold $P_{-k,n}$
$\widehat{\mu}_{-k,n}\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}P_{-k,n}[A(Y-\tilde{\mu}(X))^{2}],$
and also assumptions on the gap between the constrained and unconstrained outcome models. Also in this section, we show how these aforementioned assumptions satisfy the C-Learner assumptions required for Theorems (1) and (2).
Then in the following sections, we show how self-normalized AIPW and TMLE satisfy the assumptions stated in this section on the gap between the constrained and unconstrained outcome models.
Assumptions on unconstrained outcome models
The following assumption on unconstrained outcome models is analogous to Assumption (B). This assumption is standard.
Assumption F (Convergence rates of propensity and unconstrained outcome models).
For all $k\in[K]$,
$\|\widehat{\pi}_{-k,n}-\pi\|_{L_{2}(P)}=o_{p}(n^{-\frac{1}{4}}),\quad\|\widehat{\mu}_{-k,n}^{C}-\mu\|_{L_{2}(P)}=o_{p}(n^{-\frac{1}{4}}).$
Assumptions on the difference between constrained and unconstrained outcome models
These assumptions essentially ensure that the first-order constraint (17) does not change the outcome model too much asymptotically, i.e., $\widehat{\mu}^{C}_{-k,n}$ and $\widehat{\mu}_{-k,n}$ are similar asymptotically. This first constraint is used to show Assumption (B):
Assumption G (The constraint is negligible asymptotically).
For all $k$ = 1, …, K,
$\|\widehat{\mu}^{C}_{-k,n}-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}=o_{p}(n^{-1/4}).$
Notably, the distance between the constrained solution $\widehat{\mu}^{C}_{-k,n}$ and its unconstrained counterpart $\widehat{\mu}_{-k,n}$ can be as large as that between $\widehat{\mu}_{-k,n}$ and the true parameter $\mu$, asymptotically.
This next assumption is used to show Assumption (D):
Assumption H (Empirical process assumption on $\widehat{\mu}^{C}_{-k,n}$ vs $\widehat{\mu}_{-k,n}$).
$(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{C}(X)-\widehat{\mu}_{-k,n}(X))=o_{p}(n^{-1/2}).$
See Appendix E.1 for an example of why this assumption is necessary. Given these assumptions, we show the C-Learner assumptions hold.
Proposition 3 (C-Learner assumptions hold, given assumptions on unconstrained outcome models and the gap between constrained and unconstrained models).
Assume Assumptions (A), (C), (F), and (H). Then Assumptions (A), (C), (B), (D) are satisfied.
To show this proposition, it suffices to show Assumptions B and D. We will show these assumptions in the rest of this section.
Showing Assumption (B)
By the triangle inequality,
$\|\widehat{\mu}^{C}_{-k,n}-\mu\|_{L_{2}(P)}$ $\leq\|\widehat{\mu}_{-k,n}-\mu\|_{L_{2}(P)}+\|\widehat{\mu}^{C}_{-k,n}-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}.$ (26)
Showing Assumption (D)
By the triangle inequality,
$|T_{1k}|$ $:=\left|(P_{k,n}-P)\left(\varphi(Z;\widehat{P}_{k,n}^{C})-\varphi(Z;P)\right)\right|$
$\leq\left|(P_{k,n}-P)\left(\varphi(Z;\widehat{P}_{k,n}^{C})-\varphi(Z;\widehat{P}_{-k,n})\right)\right|+\left|(P_{k,n}-P)\left(\varphi(Z;\widehat{P}_{-k,n})-\varphi(Z;P)\right)\right|$ (27)
where $\widehat{P}_{k,n}$ is defined as
$\widehat{P}_{X;k,n}$ $:=P_{X;k,n}$
$\widehat{P}_{A\mid X;k,n}[A=1\mid X=x]$ $:=\widehat{\pi}_{-k,n}(x)$
$\widehat{P}_{Y\mid A,X;k,n}[Y\mid A=1,X=x]$ $:=\widehat{\mu}_{-k,n}(x)$
with $\widehat{\pi}_{-k,n},\widehat{\mu}_{-k,n}$ as defined in Section 6, and $\widehat{P}_{k,n}^{C}$ is defined as in Section 6. The second term in (27) is addressed using a standard argument for cross-fitting, as $P_{k,n}$ and $\widehat{P}_{-k,n}$ (and therefore $\varphi(Z;\widehat{P}_{-k,n})$) use disjoint data. In contrast, the first term is not handled by standard cross-fitting arguments: $\widehat{P}_{-k,n}$ only uses data from all but the $k$-th fold, while $\widehat{P}^{C}_{k,n}$ uses the $k$-th fold, except that $\widehat{\mu}^{C}_{-k,n}$ is made to satisfy a constraint that does use the $k$-th fold, as described in Section 6. We begin by addressing the second term, which is more standard.
Lemma 1 (Cross-fitting lemma).
Let $\widehat{f}(z)$ be a function estimated from an iid sample $Z^{N}=$ $\left(Z_{n+1},\ldots,Z_{N}\right)$, and let ${P}_{n}$ denote the empirical measure over $\left(Z_{1},\ldots,Z_{n}\right)$, which is independent of $Z^{N}$. Then (omitting arguments for brevity)
$\left({P}_{n}-{P}\right)(\widehat{f}-f)$
has mean 0 and variance that is upper bounded by
$\frac{\|\widehat{f}-f\|_{L_{2}(P)}^{2}}{n}.$
Proof    First note that the conditional mean is 0, i.e. $P\left[(P_{n}-P)(\widehat{f}-f)\mid Z^{N}\right]=0$ since
$P\left[{P}_{n}(\widehat{f}-f)\mid Z^{N}\right]=P\left(\widehat{f}-f\mid Z^{N}\right)={P}(\widehat{f}-f).$
The conditional variance is
$\operatorname{Var}\left\{\left({P}_{n}-{P}\right)(\widehat{f}-f)\mid Z^{N}\right\}=\operatorname{Var}\left\{{P}_{n}(\widehat{f}-f)\mid Z^{N}\right\}=\frac{1}{n}\operatorname{Var}\left(\widehat{f}-f\mid Z^{N}\right)\leq\frac{\|\widehat{f}-f\|_{L_{2}(P)}^{2}}{n}.$
Then for (unconditional) mean and variance, $P\left[(P_{n}-P)(\widehat{f}-f)\right]=0$ and
${\rm Var}\left\{(P_{n}-P)(\widehat{f}-f)\right\}$ $={\rm Var}\left\{P\left[(P_{n}-P)(\widehat{f}-f)\mid Z^{N}\right]\right\}+E\left[{\rm Var}\left\{(P_{n}-P)(\widehat{f}-f)\mid Z^{N}\right\}\right]$
$\leq 0+\frac{\|\widehat{f}-f\|_{L_{2}(P)}^{2}}{n}.$
Now we show the second term in the RHS in Equation (27) is $o_{P}(n^{-1/2})$. To do this, write
$\varphi(Z;P)$ $=\frac{A}{\pi(X)}(Y-\mu(X))+\mu(X)-\psi(P)$ (28)
$\varphi(Z;\widehat{P}_{-k,n})$ $=\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))+\widehat{\mu}_{-k,n}(X)-\psi(\widehat{P}_{-k,n}).$ (29)
Observe that $(P_{k,n}-P)\psi(P)=(P_{k,n}-P)\psi(\widehat{P}_{-k,n})=0$ as $\psi(P),\psi(\widehat{P}_{-k,n})$ are constants. Thus, it remains to show that for a fixed $k$, $(P_{k,n}-P)(\widehat{f}_{k,n}-f)=o_{P}(n^{-1/2})$, where we omit arguments for brevity and let
$\widehat{f}_{k,n}$ $=\frac{A}{\widehat{\pi}_{-k,n}}(Y-\widehat{\mu}_{-k,n})+\mu_{-k,n},$ (30)
$f$ $=\frac{A}{\pi}(Y-\mu)+\mu.$ (31)
We do this by using Lemma (1). Observe that
$\widehat{f}_{k,n}-f=\left(1+\frac{A}{\pi}\right)(\mu-\widehat{\mu}_{-k,n})+\frac{A}{\widehat{\pi}_{-k,n}\cdot\pi}(Y-\widehat{\mu}_{-k,n})(\pi-\widehat{\pi}_{-k,n}).$ (32)
Then using Assumption (A),
$\|\widehat{f}_{k,n}-f\|_{L_{2}(P)}$ $\leq\left(1+\frac{1}{\eta}\right)\|\widehat{\mu}_{-k,n}-\mu\|_{L_{2}(P)}+\frac{1}{\eta^{2}}\|A(Y-\widehat{\mu}_{-k,n})\|_{L_{2}(P)}\|\pi-\widehat{\pi}_{-k,n}\|_{L_{2}(P)}.$ (33)
The leftmost term on the RHS converges to 0, using Assumption (F). Note that we can bound the rightmost term using the triangle inequality
$\left(1+\frac{1}{\eta^{2}}\right)\left(\|A(Y-\mu)\|_{L_{2}(P)}+\|\mu-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}\right)\|\pi-\widehat{\pi}_{-k,n}\|_{L_{2}(P)},$ (34)
which also converges to 0 by Assumption (F) and Assumption (C). Thus combining with Lemma (1) we obtain that the second term on the RHS of Equation (27) is $o_{P}(n^{-1/2})$, as $n^{1/2}(P_{k,n}-P)(\widehat{f}_{k,n}-f)$ has mean 0 and variance $\leq\|\widehat{f}_{k,n}-f\|_{L_{2}(P)}$ which converges to 0 as $n\to\infty$.
Now we address the remaining term: we show the first term on the RHS of Equation (27) is $o_{P}(n^{-1/2})$. We use a similar argument to before: let $\widehat{f}_{k,n}$ be as before, and $\widehat{f}^{C}_{k,n}$ as below:
$\widehat{f}_{k,n}$ $:=\frac{A}{\widehat{\pi}_{-k,n}}(Y-\widehat{\mu}_{-k,n})+\widehat{\mu}_{-k,n}$ (35)
$\widehat{f}^{C}_{k,n}$ $:=\frac{A}{\widehat{\pi}_{-k,n}}(Y-\widehat{\mu}_{-k,n}^{C})+\widehat{\mu}_{-k,n}^{C}$ (36)
where we again omit the $\psi(\widehat{P}^{C}_{-k,n}),\psi(\widehat{P}_{-k,n})$ terms in $\varphi(\widehat{P}_{k,n}^{C}),\varphi(\widehat{P}_{k,n})$ from $\widehat{f}^{C}_{k,n},\widehat{f}_{k,n}$ above as they are constants. We can’t use Lemma (1) since $\widehat{f}^{C}_{k,n}$ also uses $P_{k,n}$ to fit, so instead, by using Assumptions (A) and then (H),
$(P_{k,n}-P)\left(\varphi(Z;\widehat{P}^{C}_{k,n})-\varphi(Z;\widehat{P}_{k,n})\right)$ $=(P_{k,n}-P)\left(\widehat{f}^{C}_{k,n}-\widehat{f}_{k,n}\right)$ (37)
$\leq\left(1+\frac{1}{\eta}\right)(P_{k,n}-P)(\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X))$ (38)
$=o_{P}(n^{-1/2}).$ (39)
B.4 Self-normalized AIPW satisfies Assumptions G and H
Recall that we showed how self-normalized AIPW also satisfies the C-Learner formulation in Section 3.1. Here we show that Assumptions G and H hold for self-normalized AIPW, so that Theorems 1 and 2 follow through Proposition (3).
As in the discussion in Section 3.1, self-normalized AIPW is equivalent to the specific C-Learner $\widehat{\mu}^{C}_{-k,n}$ that is defined by adjusting $\widehat{\mu}_{-k,n}$ by an additive constant:
$\widehat{\mu}^{C}_{-k,n}(x)=\widehat{\mu}_{-k,n}(x)+c_{k,n}\;\text{ where }\;c_{k,n}:=\frac{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]}{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}\right]}.$
Showing Assumption (G):
Since $\widehat{\mu}_{-k,n}^{C}$ is just a constant offset from $\widehat{\mu}_{-k,n}$, it suffices to show that $c_{k,n}=o_{P}(n^{-1/4})$, for a fixed $k$.
[…]
Now we address the numerator of $c_{k,n}$. For brevity, call this $c_{k,n}^{\text{num}}$.
$c_{k,n}^{\text{num}}$ $=P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]$ (40)
$=P\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]+(P_{k,n}-P)\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]$ (41)
$=\underbrace{P\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\mu(X))\right]}_{=0}+P\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(\mu(X)-\widehat{\mu}_{-k,n}(X))\right]$ (42)
$\quad+(P_{k,n}-P)\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]$
$=P\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(\mu(X)-\widehat{\mu}_{-k,n}(X))\right]+(P_{k,n}-P)\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]$ (43)
so that
$|c_{k,n}^{\rm num}|$ $\leq\frac{1}{\eta}P\left|\mu(X)-\widehat{\mu}_{-k,n}(X)\right|+\frac{1}{\eta}(P_{k,n}-P)\left|Y-\widehat{\mu}_{-k,n}(X)\right|$ (44)
[…]
$\|Y-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}$ $\leq\|Y-\mu\|_{L_{2}(P)}+\|A(\mu-\widehat{\mu}_{-k,n})\|_{L_{2}(P)}$
$\leq B+\|\mu-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}$
$=B+o_{P}(n^{-1/4}).$
We arrive at the desired result $c_{k,n}=o_{P}(n^{-1/4})$ as $c_{k,n}=c_{k,n}^{\rm num}/c_{k,n}^{\rm den}$, and we just showed that $1/c_{k,n}^{\rm den}=O_{P}(1)$ and $c_{k,n}^{\rm num}=o_{P}(n^{-1/4})$. Note that we have shown Assumption (G) for any sequence of $\widehat{\pi}_{-k,n}$’s, as they are bounded by a constant.
Showing Assumption (H):
To show $(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{C}(X)-\widehat{\mu}_{-k,n}(X))=o_{P}(n^{-1/2})$, in the case of the constant shift $\widehat{\mu}^{C}(X)=\widehat{\mu}(X)+c_{k,n}$, note that $\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X)=c_{k,n}$ for every $X$. Therefore,
$(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{C}(X)-\widehat{\mu}_{-k,n}(X))=(P_{k,n}-P)c_{k,n}=0,$
with $c_{k,n}<\infty$ a.s. for large enough $n$ (which exists by Borel-Cantelli, as in when we showed Assumption (G)).
B.5 TMLE satisfies Assumptions G and H
Recall that we showed how a version of the TMLE for estimating the ATE with continuous unbounded outcomes also satisfies the C-Learner formulation in Section 3.1. Here we show that a cross-fitted version of the TMLE for estimating the ATE with continuous unbounded outcomes additionally satisfies Assumptions (G) and (H), so that Theorems (1) and 2 follow through Proposition (3). Note that the formulation below fits a separate $\epsilon^{\star}_{k,n}$ per cross-fitting split for consistency with C-Learner, rather than one $\epsilon^{\star}$ overall as described in the cross-validated version of TMLE in [40].
$\widehat{\mu}^{C}_{-k,n}(X)=\widehat{\mu}_{-k,n}(X)+\epsilon^{\star}_{k,n}\frac{A}{\widehat{\pi}_{-k,n}(X)},\;\text{ where }\epsilon^{\star}_{k,n}=\frac{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]}{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}^{2}(X)}\right]}.$
Showing Assumption (G):
Observe that by Assumption (A)
$\left\|\epsilon_{k,n}^{\star}\frac{A}{\widehat{\pi}_{-k,n}(X)}\right\|_{L_{2}(P)}\leq\frac{1}{\eta}\|\epsilon_{k,n}^{\star}\|_{L_{2}(P)}$
so it suffices to show that $\epsilon_{k,n}^{\star}=o_{P}(n^{-1/4})$. Note that since $\widehat{\pi}_{-k,n}(x)\leq 1$ for all $x$,
$|\epsilon_{k,n}^{\star}|\leq\frac{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\widehat{\mu}_{-k,n}(X))\right]}{P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}\right]}\leq|c_{k,n}|$ (45)
where $c_{k,n}$ is the constant adjustment in Section B.4, so that $\epsilon^{\star}_{k,n}=o_{P}(n^{-1/4})$.
Showing Assumption (H):
We want to show
$(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{\star}(X)-\widehat{\mu}_{-k,n}(X))=(P_{k,n}-P)\left(\epsilon^{\star}_{k,n}\frac{A}{\widehat{\pi}_{-k,n}(X)}\right)=o_{P}(n^{-1/2}).$
Note that
$(P_{k,n}-P)\left(\epsilon^{\star}_{k,n}\frac{A}{\widehat{\pi}_{-k,n}(X)}\right)=\epsilon^{\star}_{k,n}(P_{k,n}-P)\left(\frac{A}{\widehat{\pi}_{-k,n}(X)}\right)$
and that $(P_{k,n}-P)\left(\frac{A}{\widehat{\pi}_{-k,n}(X)}\right)=O_{P}(n^{-1/2})$ by Lemma (1), since $|A/\widehat{\pi}_{-k,n}(X)|\leq 1/\eta$ a.s. by Assumption (A). Additionally, $\epsilon^{\star}_{k,n}=o_{P}(n^{-1/4})$, by the argument above where we showed Assumption (G). The desired result follows by taking the product of the rates for $\epsilon^{\star}_{k,n}$ and $(P_{k,n}-P)\left(\frac{A}{\widehat{\pi}_{-k,n}(X)}\right)$.
Appendix C Additional experiment details and results
C.1 Kang & Schafer experiment
C.1.1 Additional Kang & Schafer experiment details
Linear outcome models.
Here, $P_{\rm train}=P_{\rm eval}$, as consistent with the original paper by Kang and Schafer [23] [^2] Although sample splitting could be better, we have chosen to replicate [23] as closely as possible.. There is no $P_{\rm val}$ as there are no hyperparameters to tune.
For the direct method and initial outcome model for AIPW, self-normalized AIPW, and TMLE, we simply regress the dependent variable $Y$ on the observed covariates $X$ using the samples with labels. We use all the samples available to fit a logistic regression model for the treatment variable $A$ (outcome was observed) using the covariates $X$. Following [23], in both models, we omit the intercept term when fitting the propensity scores and initial outcome model. Our main insights do not change if an intercept is added to the outcome and propensity models.
We run datasets with 200 and 1000 observations, with and without a clip of $5\%$. For each configuration, we generate 1000 seeds.
Table 4 displays the results for the bias, mean absolute error, root mean squared error (RMSE), and median absolute error over 1000 simulations with sample sizes equal to 200 and 1000. We also include the results when clipping of 5% is performed, meaning that $\widehat{\pi}(X)$ is truncated to have a lower bound of 0.05. Note that clipping greatly improves the performance of propensity-based methods such as AIPW and TMLE. Without clipping, C-Learner is the best method across all samples when compared to other unbiased methods and is outperformed only by the direct method. When clipping is introduced, the C-Learner is similar in performance to other asymptotically optimal methods.
For the coverage computation, we construct the confidence intervals as described in Appendix D and empirically check for each method if it covers the true populational mean. Table 6 presents the coverage results for a 95% confidence level. It is worth noticing that coverage results greatly deteriorate when the sample size increases for all asymptotically optimal methods, most likely because both the outcome model and the propensity model are mispecified, and consistency is ensured when at least one of the models is consistent (see Section 6).
Finally, Table 7 displays the results for bias and mean absolute error (MAE) for estimating $P[Y(0)]$ rather than $P[Y(1)]$. In this case, all the asymptotic optimal methods perform very similarly, and all are better than the direct method with OLS. To conclude, we emphasize that the C-Learner is the only asymptotically optimal method that demonstrated good performance across all configurations of data size, clipping procedure, and data-generating process by either achieving comparable performance to other asymptotically optimal methods, or by improving performance, sometimes by an order of magnitude.
C.1.2 Additional Kang & Schafer results
We present in Table 4 the mean absolute error, root mean square error and median absolute error for the linear specification described in Section 5.1 with their respective standard errors for different sample sizes 5% clipping. We highlight the best performance estimator other than the direct method.
(a) $N=200$, no clip
method Bias Mean Abs Err RMSE Median Abs Err
Direct -0.00 (0.103) 2.60 (0.061) 3.24 (0.457) 2.13
IPW 22.10 (2.579) 27.28 (2.52) 84.45 (2733) 9.87
IPW-POP 3.36 (0.292) 5.42 (0.260) 9.83 (13.45) 3.01
AIPW -5.08 (0.474) 6.16 (0.461) 15.82 (87.1) 3.43
AIPW-POP -3.65 (0.203) 4.73 (0.179) 7.38 (6.14) 3.26
TMLE -111.59 (41.073) 112.15 (41.1) 1302.98 ($10^{6}$) 3.95
C-Learner -2.45 (0.120) 3.57 (0.088) 4.52 (0.912) 2.93
(b) $N=1000$, no clip
Method Bias Mean Abs Err RMSE Median Abs Err
Direct -0.43 (0.044 ) 1.17 (0.028) 1.46 (0.095) 1.00
IPW 105.46 (59.843) 105.67 (59.8) 1894.40 ($10^{6}$) 17.95
IPW-POP 6.83 (0.331) 7.02 (0.327) 12.50 (24.0) 4.09
AIPW -41.37 (24.821) 41.39 (24.8) 785.60 ($10^{5}$) 5.22
AIPW-POP -8.35 (0.431) 8.37 (0.430) 15.97 (46.4) 4.92
TMLE -17.51 (3.493) 17.51 (3.49) 111.77 ($10^{4}$) 4.25
C-Learner -4.40 (0.077) 4.42 (0.076) 5.03 (0.795) 4.21
(c) $N=200$, clip 5%
Method Bias Mean Abs Err RMSE Median Abs Err
Direct -0.00 (0.103) 2.60 (0.061) 3.24 (0.457) 2.13
IPW 5.13 (0.398) 10.47 (0.275) 13.60 (10.055) 8.35
IPW-POP 1.17 (0.132) 3.33 (0.087) 4.32 (0.969) 2.59
AIPW -2.45 (0.121) 3.59 (0.088) 4.54 (0.913) 3.02
AIPW-POP -2.37 (0.118) 3.50 (0.085) 4.42 (0.859) 2.92
TMLE -2.01 (0.107) 3.15 (0.075) 3.95 (0.665) 2.62
C-Learner -2.15 (0.112) 3.31 (0.079) 4.14 (0.734) 2.83
(d) $N=1000$, clip 5%
Method Bias Mean Abs Err RMSE Median Abs Err
Direct -0.43 (0.044) 1.17 (0.028) 1.46 (0.095) 1.00
IPW 8.35 (0.179) 8.67 (0.163) 10.08 (3.413) 8.23
IPW-POP 1.95 (0.062) 2.27 (0.050) 2.76 (0.293) 2.01
AIPW -3.41 (0.053) 3.44 (0.051) 3.80 (0.378) 3.44
AIPW-POP -3.31 (0.052) 3.34 (0.050) 3.70 (0.358) 3.35
TMLE -2.81 (0.046) 2.85 (0.044) 3.17 (0.274) 2.86
C-Learner -3.07 (0.049) 3.10 (0.047) 3.43 (0.310) 3.10
Table 4: Estimator performance in 1000 tabular simulations for the linear specification of outcome models. “Clip” refers to clipping $\widehat{\pi}(X)$ away from 0. Standard error is in parentheses. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing overall method other than the direct method in bold.
(a) $N=200$, clip = 5%
Method Bias Mean Abs Err
Direct -6.10 (0.10) 6.18 (0.10)
IPW 4.45 (0.75) 17.8 (0.52)
IPW-POP -3.40 (0.16) 5.06 (0.10)
Lagrangian -4.94 (0.10) 5.14 (0.09)
AIPW -1.80 (0.14) 3.85 (0.09)
AIPW-POP -2.10 (0.12) 3.64 (0.08)
TMLE -2.84 (0.10) 3.51 (0.08)
C-Learner -3.10 (0.10) 3.68 (0.08)
(b) $N=200$, clip = 0.1%
Bias Mean Abs Err
-6.10 (0.10) 6.18 (0.10)
41 (5.82) 54.2 (5.72)
-1.10 (0.38) 7.20 (0.31)
-4.96 (0.10) 5.16 (0.09)
4.82 (1.27) 10.4 (1.23)
-0.57 (0.26) 5.02 (0.21)
-1.42 (0.17) 4.19 (0.12)
-3.24 (0.09) 3.79 (0.08)
Table 5: Comparison of estimator performance on misspecified datasets from Kang and Schafer [23] in 1000 tabular simulations using gradient boosted regression trees with 200 samples, 5%, and 0.1% clipping. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing method in bold. Standard errors are displayed within parentheses to the right of the point estimate. “Lagrangian” refers to only performing the first stage in Algorithm 1.
$N=200$ $N=1000$
Method No Clip 5% Clip No Clip 5% Clip
Direct 0.89 0.89 0.88 0.88
IPW 1.00 1.00 1.00 1.00
IPW-POP 1.00 1.00 1.00 1.00
AIPW 0.88 0.90 0.59 0.56
AIPW-POP 0.92 0.91 0.72 0.58
TMLE 0.84 0.90 0.48 0.60
C-Learner 0.91 0.90 0.68 0.58
Table 6: Coverage results for 1000 simulations in the linear specification. Confidence intervals were set to achieve 95% confidence. Asymptotically optimal methods are listed beneath the horizontal divider.
(a) $N=200$
Method Bias Mean Abs Err
Direct 4.82 (0.091) 4.94 (0.085)
IPW -0.70 (0.141) 3.44 (0.092)
IPW-POP 2.48 (0.097) 3.23 (0.072)
AIPW 3.23 (0.093) 3.63 (0.077)
AIPW-POP 3.22 (0.092) 3.61 (0.076)
TMLE 3.11 (0.091) 3.50 (0.076)
C-Learner 3.19 (0.092) 3.58 (0.076)
(b) $N=1000$
Bias Mean Abs Err
4.80 (0.041) 4.80 (0.041)
-0.83 (0.054) 1.52 (0.036)
2.43 (0.042) 2.47 (0.040)
3.14 (0.041) 3.15 (0.040)
3.12 (0.041) 3.13 (0.040)
3.03 (0.041) 3.04 (0.040)
3.09 (0.041) 3.10 (0.040)
Table 7: Results of 1000 simulations for estimating $P[Y(0)]$ instead of $P[Y(1)]$ with the linear specification. The standard error is in parentheses. Asymptotically optimal methods are listed beneath the horizontal divider.
Method $N=200$ $N=1000$
Direct 0.35 (0.01) 0.12 (0.01)
IPW 1.00 (0.00) 0.98 (0.01)
IPW-POP 0.85 (0.01) 1.00 (0.00)
Lagrangian 0.84 (0.01) 0.71 (0.01)
AIPW 0.85 (0.01) 0.85 (0.01)
AIPW-POP 0.85 (0.01) 0.85 (0.01)
TMLE 0.86 (0.01) 0.87 (0.01)
C-Learner 0.82 (0.01) 0.77 (0.01)
Table 8: Coverage results for 1000 simulations using gradient boosted regression trees. Confidence intervals were set to achieve 95% confidence. Asymptotically optimal methods are listed beneath the horizontal divider. “Lagrangian” refers to only performing the first stage in Algorithm 1.
Gradient boosted regression trees.
We demonstrate the flexibility and performance of the C-Learner in which we instantiate C-Learner using gradient boosted regression trees using the XGBoost package [9] with a custom objective, as outlined in Section 4.2. $\widehat{\pi}$ is fit as a logistic regression on covariates $X$ using L1 regularization (LASSO).
For each seed, and sample size, we randomly take half of the data for $P_{\rm train}$ and half of the data for $P_{\rm eval}$. For the first phase of Algorithm 1, we perform hyperparameter tuning using $P_{\rm val}=P_{\rm eval}$. Hyperparameter tuning is performed using a grid search for the following parameters: learning rate (0.01, 0.05, 0.1, 0.2), feature subsample by tree (0.5, 0.8, 1), and max tree depth (3, 4, 5). We also set the maximum number of weak learners to 2 thousand, and we perform early stopping using MSE loss on $P_{\rm val}$ for 20 rounds. Hyperparameter tuning is performed separately for C-Learner and the initial outcome model.
For the second phase of Algorithm 1, we use the set of hyperparameters found in the first stage. The weak learners are fitted using $P_{\rm eval}$ and $P_{\rm val}=P_{\rm eval}$. In order to avoid overfitting in the targeting step, we use a subsampling of 50% and early stopping after 20 rounds.
In Table 5 we provided additional details with clipping in order to understand the impact in the results of AIPW and TMLE that produced unreliable estimates for sample sizes of 200 and 1000 as well as the impact of clipping in the other estimates. When “small” clipping is performed (0.1%), C-Learner is still better than any other estimator. When clipping of 5% is performed, TMLE achieves the best pointwise performance, and C-Leaner and AIPW-POP are statically equal in terms of MAE.
Finally, in Table 8 we present the coverage results for the estimators computed without clipping and sample size of 200 and 1000 presented in the main text on Table  2. For a sample size of 200, asymptotically optimal methods have similar coverage and substantially lower than the target of 95%. When the sample size increases, the coverage results of C-Learner deteriorate, although it still presents the best performance in terms of MAE.
C.2 CivilComments experiment
C.2.1 Additional CivilComments experiment details
Model and training objectives
Both propensity and outcome models are a linear layer on top of a DistilBERT featurizer, with either softmax and cross-entropy loss for propensity score, or mean squared error for outcome models. For training outcome models, we only consider training data on which $A=1$, as only those terms contribute to the loss.
Hyperparameters and settings
For propensity models, we tried learning rates of $\{10^{-3},10^{-4},10^{-5}\}$ for the setting of $l=10^{-4}$ and found that a learning rate of $10^{-4}$ performed the best in terms of val loss. Because of computational constraints, we also used this learning rate for $l=10^{-2},10^{-3}$.
For outcome models (including C-Learner) in the settings of $l\in\{10^{-2},10^{-3},10^{-4}\}$, we tried learning rates of $\{10^{-3},10^{-4},10^{-5},10^{-6}\}$. For C-Learner regularization, we tried $\lambda$ taking on values of $\lambda_{0}/P_{\rm eval}[A/\hat{\pi}(X)]^{2}$ where $\lambda_{0}\in\{0,1,4,16,64,256\}$. Essentially, the penalty is $\lambda_{0}$ times the square of the bias shift that would be required to satisfy the condition.
Because of computational constraints, for choosing these hyperparameters (learning rate, $\lambda$), we select a subset of dataset draws, and choose the best hyperparameters based on each model’s criteria on that small set of dataset draws. Then, after the hyperparameters are chosen, we run over the entire set of dataset draws.
While hyperparameters for the (usual) outcome model are chosen to minimize val loss, in contrast, the hyperparameters (learning rate, regularization $\lambda$) for the C-Learner outcome model were selected to minimize the magnitude of the constant shift at the end of the first epoch, as in Section 4.3. The idea is that if the size of the constant shift is small, then the regularizer is doing a good job of enforcing the constraint. Although one might think that larger regularization $\lambda$’s would automatically lead to smaller constant shifts, we do not find this to be the case; the best $\lambda$ in our settings (64) is not the largest value that we try (up to 256). Furthermore, the set of hyperparameter values we choose between in this process are ones that seem to result in reasonable performance, e.g. in terms of MSE on treated units in the validation set.
The best hyperparameters chosen are in Table 9. We trained all models with batch size 64. Learning rates decayed linearly over 10 epochs. Optimization was done using the AdamW optimizer. Weight decay was fixed at $0.1$.
We choose the training epoch with the best val loss for each model (cross entropy for propensity, and MSE for outcome). For C-Learner, we choose the epoch with the best val MSE.
C.2.2 Additional CivilComments results
We consider three data generating processes as in Section 5.2, where treatment assignment is parameterized by $l=10^{-4},10^{-3},10^{-2}$, respectively. Hyperparameters for outcome models are in Table 9. Results are in Tables 101112.
$l=10^{-4}$ $l=10^{-3}$ $l=10^{-2}$
Outcome model learning rate $10^{-3}$ $10^{-3}$ $10^{-4}$
C-Learner learning rate (best val MSE) $10^{-4}$ $10^{-4}$ $10^{-4}$
C-Learner learning rate (min bias shift) $10^{-5}$ $10^{-5}$ $10^{-5}$
C-Learner $\lambda$ (best val MSE) 0 16 16
C-Learner $\lambda$ (min bias shift) 64 64 64
Table 9: Hyperparameters for $l=10^{-4},10^{-3},10^{-2}$ for Section 5.2
Method Bias Mean Abs Err Coverage
Direct 0.173 (0.008) 0.177 (0.007) 0.010 (0.001)
IPW 0.504 (0.084) 0.546 (0.081) 0.760 (0.018)
IPW-POP 0.114 (0.017) 0.153 (0.014) 0.890 (0.010)
AIPW 0.084 (0.043) 0.307 (0.032) 0.830 (0.014)
AIPW-POP 0.116 (0.018) 0.161 (0.014) 0.850 (0.013)
TMLE -1.264 (1.361) 1.802 (1.355) 0.810 (0.015)
C-Learner (best val MSE) 0.103 (0.015) 0.141 (0.011) 0.870 (0.011)
C-Learner (min bias shift) 0.075 (0.012) 0.115 (0.008) 0.900 (0.009)
Table 10: Comparison of estimators in the CivilComments [13] semi-synthetic dataset over 100 re-drawn datasets, with $l=10^{-4}$. Confidence intervals were set to achieve 95% confidence. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing overall method in bold. Standard errors are displayed within parentheses to the right of the point estimate.
Method Bias Mean Abs Err Coverage
Direct 0.147 (0.028) 0.200 (0.025) 0.000 (0.000)
IPW 0.417 (0.064) 0.437 (0.063) 0.840 (0.013)
IPW-POP 0.079 (0.015) 0.120 (0.013) 0.910 (0.008)
AIPW -0.056 (0.052) 0.344 (0.039) 0.950 (0.005)
AIPW-POP 0.089 (0.020) 0.134 (0.018) 0.950 (0.005)
TMLE 0.074 (0.025) 0.144 (0.022) 0.940 (0.006)
C-Learner (best val MSE) 0.067 (0.012) 0.110 (0.009) 0.980 (0.002)
C-Learner (min bias shift) 0.060 (0.011) 0.099 (0.008) 0.990 (0.001)
Table 11: Comparison of estimators in the CivilComments [13] semi-synthetic dataset over 100 re-drawn datasets, with $l=10^{-3}$. Confidence intervals were set to achieve 95% confidence. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing overall method in bold. Standard errors are displayed within parentheses to the right of the point estimate.
Method Bias Mean Abs Err Coverage
Direct 0.091 (0.007) 0.099 (0.006) 0.040 (0.004)
IPW 0.346 (0.041) 0.353 (0.040) 0.710 (0.021)
IPW-POP 0.004 (0.005) 0.038 (0.003) 0.950 (0.005)
AIPW -0.172 (0.029) 0.220 (0.026) 0.890 (0.010)
AIPW-POP 0.002 (0.005) 0.040 (0.003) 1.000 (0.000)
TMLE 0.027 (0.005) 0.042 (0.003) 0.990 (0.001)
C-Learner (best val MSE) 0.009 (0.007) 0.053 (0.004) 1.000 (0.000)
C-Learner (min bias shift) 0.001 (0.006) 0.045 (0.003) 1.000 (0.000)
Table 12: Comparison of estimators in the CivilComments [13] semi-synthetic dataset over 100 re-drawn datasets, with $l=10^{-2}$. Confidence intervals were set to achieve 95% confidence. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing asymptotically optimal method in bold. Standard errors are displayed within parentheses to the right of the point estimate. C-Learner’s performance is within standard error of AIPW-POP and TMLE.
C-Learner errors are comparatively more stable for $l=10^{-4},10^{-3}$. For $l=10^{-2}$, C-Learner is comparable to other asymptotically optimal methods.
C.3 IHDP Dataset (tabular)
The Infant Health and Development Program (IHDP) dataset is a tabular semisynthetic dataset [6] that was first introduced as a benchmark for ATE estimation by Hill [20]. IHDP is based on a randomized experiment studying the effect of specialized healthcare interventions on the cognitive scores of premature infants with low birth weight. The dataset consists of a continuous outcome variable $Y$, a binary treatment variable $A$, and 25 covariates $X$ that affect the outcome variable and are also correlated with the treatment assignment. We use 1000 datasets as in [38, 12].
C.3.1 Gradient boosted regression trees
We instantiate C-Learner using gradient boosted regression trees using the XGBoost package [9] with a custom objective, as outlined in Section 4.2. $\widehat{\pi}$ is fit as a logistic regression on covariates $X$. Different to the applications presented in the main text, we are interested in estimating the ATE $P[Y(1)-Y(0)]$ where $Y(0)$ is nonzero. We discuss two alternatives to instantiate the C-Learner. First, we find an outcome model that takes as input both covariates and the treatment variable, which leads to the following formulation
$\widehat{\mu}^{C}_{-k,n}=\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{-k,n}[(Y-\tilde{\mu}(A,X))^{2}]:P_{k,n}\left[\left(\frac{A}{\widehat{\pi}(X)}-\frac{1-A}{1-\widehat{\pi}(X)}\right)(Y-\tilde{\mu}(A,X))\right]=0\right\},$
and the final estimator becomes
$\widehat{\psi}^{C}_{n}:=\frac{1}{K}\sum_{k=1}^{K}P_{k,n}[\widehat{\mu}_{-k,n}^{C}(1,X)-\widehat{\mu}_{-k,n}^{C}(0,X)].$
This approach is often referred to as the “S-Learner”.
Another alternative is to model the ATE as the difference of two missing outcomes and fit two outcome models for treated vs. non-treated units. This approach is commonly referred to as the “T-Learner”. In this case, we estimate two models by solving
$\widehat{\mu}_{-k,n,1}^{C}$ $\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{-k,n}[A(Y-\tilde{\mu}(X))^{2}]:P_{k,n}\left[\frac{A}{\widehat{\pi}_{-k,n}(X)}(Y-\tilde{\mu}(X))\right]=0\right\},$
$\widehat{\mu}_{-k,n,0}^{C}$ $\in\mathop{\rm argmin}_{\tilde{\mu}\in\mathcal{F}}\left\{P_{-k,n}[(1-A)(Y-\tilde{\mu}(X))^{2}]:P_{k,n}\left[\frac{1-A}{1-\widehat{\pi}_{-k,n}(X)}(Y-\tilde{\mu}(X))\right]=0\right\},$
and the final estimator becomes
$\widehat{\psi}^{C}_{n}:=P_{\rm eval}\frac{1}{K}\sum_{k=1}^{K}P_{k,n}[\widehat{\mu}_{-k,n,1}^{C}(X)-\widehat{\mu}_{-k,n,0}^{C}(X)].$
We empirically observe that the latter approach (T-Learner) is critical for performance in this setting: using a simple XGBoost T-Learner model significantly improves upon sophisticated S-Learner variants, including state-of-the-art methods such as RieszNet, RieszForest, and DragonNet [12, 38].
For consistency with prior work  [41, 12, 38], for each dataset run, we split the data in 80% training and 20% for hyperparameter tuning, and use the whole dataset when evaluating the first-order error term, i.e., $P_{\rm eval}=P_{\rm train}\cup P_{\rm val}$. For each outcome model, $\widehat{\mu}_{-k,n,1}^{C}$ and $\widehat{\mu}_{-k,n,0}^{C}$, we follow the same setup described in the Kang & Schafer experiment. Hyperparameter tuning is performed using a grid search for the following parameters: learning rate (0.01, 0.05, 0.1, 0.2), feature subsample by tree (0.5, 0.8, 1), and max tree depth (3, 4, 5). We also set the maximum number of weak learners to 2000, and we perform early stopping using MSE loss on $P_{\rm val}$ for 20 rounds. Hyperparameter tuning is performed separately for C-Learner and the initial outcome model. For the second phase of Algorithm  1, we use the set of hyperparameters found in the first stage. The weak learners are fitted using $P_{\rm eval}$. In order to avoid overfitting in the targeting step, we use a subsampling of 50% and early stopping after 20 rounds.
The results for the mean absolute error and their respective standard errors are displayed in Table 13. The C-Learner improves upon the direct method while using an identical model class. The asymptotically optimal estimators perform similarly, and C-Learner’s advantage over other asymptotically optimal methods is not statistically significant. This is perhaps not surprising in this setting as the propensity weights do not vary as much here, in contrast to Kang & Shafer’s example where existing asymptotically optimal methods performed poorly as estimated propensity weights varied dramatically.
Method Bias Mean Abs Err
Direct (Boosting) -0.16 (0.004) 0.188 (0.004)
IPW -1.50 (0.040) 1.500 (0.040)
IPW-POP -0.01 (0.003) 0.110 (0.003)
Lagrangian -0.09 (0.004) 0.144 (0.004)
AIPW -0.05 (0.002) 0.103 (0.003)
AIPW-POP -0.04 (0.002) 0.103 (0.003)
TMLE 0.007 (0.003) 0.103 (0.003)
C-Learner -0.004 (0.003) 0.104 (0.003)
Table 13: Comparison of gradient boosted regression tree-based estimators in the IHDP semi-synthetic dataset over 1000 simulations. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing method in bold. Standard errors are displayed within parentheses to the right of the point estimate. “Lagrangian” refers to only performing the first stage in Algorithm 1. C-Learner’s performance is within standard error of AIPW-POP.
C.3.2 Neural networks
We also instantiate C-Learner as neural networks (Section 4.3 on the IHDP dataset. In particular, we demonstrate how C-Learner can use Riesz representers (A), or equivalently, propensity models in the ATE setting, that are learned by other methods. This demonstrates the versatility of C-Learner as it is able to leverage new methods for learning Riesz representers. In Table 14, we demonstrate C-Learner using Riesz representers learned by RieszNet [12]. C-Learner achieves very similar performance to RieszNet while using the same Riesz representer. All methods use an outcome model trained in the usual way, except for RieszNet and C-Learner. In all settings where applicable, we use the Riesz representer ($A/\pi(X)-(1-A)/(1-\pi(X))$ learned by RieszNet.
Method Bias Mean Abs Err Coverage
Direct -0.005 (-0.000) 0.118 (0.003) 0.783 (0.025)
IPW -0.789 (-0.025) 0.903 (0.034) 0.455 (0.014)
IPW-POP -0.449 (-0.014) 0.654 (0.035) 0.819 (0.026)
AIPW -0.044 (-0.001) 0.106 (0.003) 0.940 (0.030)
AIPW-POP -0.042 (-0.001) 0.106 (0.003) 0.961 (0.030)
RieszNet (“DR”) -0.033 (-0.001) 0.098 (0.003) 0.972 (0.031)
C-Learner -0.038 (-0.001) 0.098 (0.002) 0.955 (0.030)
Table 14: Comparison of neural network based estimators in the IHDP semi-synthetic dataset over 1000 simulations. Asymptotically optimal methods are listed beneath the horizontal divider. We highlight the best-performing asymptotically optimal method in bold. Standard errors are displayed within parentheses to the right of the point estimate. All methods use an outcome model trained in the usual way, except for RieszNet and C-Learner. In all settings where applicable, we use the Riesz representer learned by RieszNet.
Training procedure
We use lightly modified RieszNet code [12] to generate the Riesz representers used in all methods. We use the same neural network architecture for RieszNet outcome models for C-Learner. We use $P_{\rm train},P_{\rm val},P_{\rm eval}$ as in Appendix C.3.1. Hyperparameters (learning rate, $\lambda$) are selected for best mean squared error on $P_{\rm val}$ for each individual dataset. For the training objective, we largely follow Section 4.3, with the following modifications for the full ATE rather than assuming that $Y(0)=0$: we use the mean squared error across the full dataset (as it was originally for RieszNet), and we have separate penalties and bias shifts for $A=1$ and $A=0$. As in RieszNet training, there is a phase of “pre-training” with a higher learning rate, and then a phase with a lower learning rate; for “pre-training” we sweep over learning rates of $\{10^{-3},10^{-4}\}$ for 100 epochs; regular training uses a learning rate of $10^{-5}$ for 600 epochs. Both use the Adam optimizer. For $\lambda$ we sweep over values of $\{0,0.01,0.02,0.04,0.1,0.2,0.4,1,2,4,8\}$. Our implementation of RieszNet as a baseline is almost identical to that of the original paper, except for how we set the random seed, for initializing neural networks and for choosing train and test split.
Comparisons with RieszNet paper
Surprisingly, our mean absolute error results in Section C.3.2, including the ones using regular RieszNet (including their algorithm code and hyperparameters), often outperform the results reported in the original RieszNet paper [12]. One difference between our code and the RieszNet code is how random seeds were set. The random seed affects the train/test split, and also weight initialization for the neural network. We re-ran the original RieszNet code, and then again with a fixed seed (set to 123) for every dataset, and we found that the results were noticeably different. In particular, the RieszNet mean absolute error is smaller with seeds set the second way. Compare Table 15 and Table 16. In our experiments on this setting, we fix seeds in the second way.
Method Mean absolute error (s.e.)
Doubly robust 0.115 (0.003)
Direct 0.124 (0.004)
IPW 0.801 (0.040)
Table 15: Original RieszNet IHDP results
Method Mean absolute error (s.e.)
Doubly robust 0.098 (0.003)
Direct 0.118 (0.003)
IPW 0.903 (0.034)
Table 16: RieszNet IHDP results, with fixed seeds
Relative performance
We’ve seen in Section 5.2 that C-Learner tends to outperform one-step estimation and targeting in settings with low overlap. Here, it merely performs about the same as one-step estimation and targeting. It seems as though overlap is not very low in IHDP, however, based on propensity scores from trained models not taking on extreme values. Using RieszNet’s learned Riesz representers, the smallest value for $\textrm{min}(\widehat{\pi},1-\widehat{\pi})$ across all IHDP datasets is 0.11, which is a fair bit of overlap, especially in comparison to, say, the settings with low overlap in Section 5.2 where C-Learner outperformed one-step estimation and targeting methods. To recap, over all of our experiments, C-Learner appears to perform either comparably to or better than one-step estimation and targeting, and outperform one-step estimation and targeting in settings with less overlap.
Appendix D ATE estimators and confidence interval calculations
We calculate asymptotic variance for estimators considered in the experiments, which are then used to calculate confidence intervals and coverage in the experiments.
D.1 Direct Method
The direct method involves the outcome model fitted on the partition $P_{-k,n}$ and evaluated on $P_{k,n}$. The estimator is given by:
$\widehat{\psi}_{n}^{\text{direct}}=\frac{1}{K}\sum_{k=1}^{K}\widehat{\psi}_{k,n}^{\text{direct}}=\frac{1}{K}\sum_{k=1}^{K}P_{k,n}[\widehat{\mu}_{-k,n}(X)].$ (46)
For clarity, albeit with a slight abuse of notation, let $n$ denote the number of data points in the $k$-th fold of the evaluation set. Denote by ${\rm Var}_{k,n}(X)=P_{k,n}(X-P_{k,n}X)^{2}$ the sample variance of $X$. The variance of the estimator $\widehat{\psi}_{k,n}^{\text{direct}}$ evaluated on $k$-th fold is then
${\rm Var}(\widehat{\psi}_{k,n}^{\text{direct}})=\frac{1}{n}{\rm Var}(\widehat{\mu}_{-k,n}(X))\asymp\frac{1}{n}{\rm Var}_{k,n}(\widehat{\mu}_{-k,n}(X)),$
from which we estimate the variance of the direct-method estimator.
D.2 Inverse Propensity Weighting (IPW)
Under similar notations, the inverse probability weighting method gives the estimator
$\widehat{\psi}_{n}^{\text{IPW}}=\frac{1}{K}\sum_{k=1}^{K}\widehat{\psi}_{k,n}^{\text{IPW}}=\frac{1}{K}\sum_{k=1}^{K}P_{k,n}\left[\frac{AY}{\widehat{\pi}_{-k,n}(X)}\right].$
The variance of the estimator is given by
${\rm Var}(\widehat{\psi}_{k,n}^{\text{IPW}})=\frac{1}{n}{\rm Var}\left(\frac{AY}{\widehat{\pi}_{-k,n}(X)}\right)\asymp\frac{1}{n}{\rm Var}_{k,n}\left(\frac{AY}{\widehat{\pi}_{-k,n}(X)}\right).$
D.3 Self-Normalized IPW (IPW-POP, or Hajek)
Consider $\widehat{\pi}$. Then the Hajek (self-normalized IPW) estimator is
$\widehat{\psi}_{\text{Hajek}}=\frac{\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}Y_{i}}{\widehat{\pi}(X_{i})}}{\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}}{\widehat{\pi}(X_{i})}}$
Let $\tilde{\pi}(X_{i})=\widehat{\pi}(X)\cdot\frac{1}{n}\sum\frac{A_{i}}{\widehat{\pi}(X_{i})}$ so that
$\widehat{\psi}_{\text{Hajek}}=\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}Y_{i}}{\tilde{\pi}(X_{i})}$
If we assume that $\tilde{\pi}(X_{i})$ does not change with $n$ (a false assumption, but a reasonable simplification in this case), then we can consider $\frac{A_{i}Y_{i}}{\tilde{\pi}(X_{i})}$ to be IID, and then we can apply the CLT. Then we have
$\sqrt{n}(\widehat{\psi}_{\text{Hajek}}-\psi)=N\left(0,{\rm Var}\left(\frac{AY}{\tilde{\pi}(X)}\right)\right).$
D.4 Any asymptotically optimal method
This applies to the AIPW, normalized AIPW, TMLE and C-Learner. These are asymptotically optimal and have asymptotic variance ${\rm Var}(\varphi(Z;P))$, where $\varphi$ is given in (20). Therefore, we use the plug-in for $\varphi(Z;P)$ and take its empirical variance.
Appendix E Additional proof notes
E.1 Why we have Assumption H
In settings where $\widehat{\mu}^{C}_{-k,n}$ is independent of $P_{k,n}$, Assumption G can be used to show Assumption H, e.g using Lemma 1. As a reminder, these assumptions are as follows:
  • Repeat of Assumption G: For all $k$,
    $\|\widehat{\mu}^{C}_{-k,n}-\widehat{\mu}_{-k,n}\|_{L_{2}(P)}=o_{P}(n^{-1/4}).$
  • Repeat of Assumption H:
    $(P_{k,n}-P)(\widehat{\mu}_{-k,n}^{C}(X)-\widehat{\mu}_{-k,n}(X))=o_{P}(n^{-1/2}).$
However, Assumption G may not imply Assumption H in our setting, as $\widehat{\mu}^{C}_{-k,n}$ is constructed to depend on $P_{k,n}$. If we allow $\widehat{\mu}^{C}_{-k,n}$ to be related to $\widehat{\mu}_{-k,n}$ in an arbitrary way that follows Assumption G, we can construct $\widehat{\mu}^{C}_{-k,n},\widehat{\mu}_{-k,n}$ where Assumption H does not hold. Here is an example:
Let $\widehat{\mu}_{-k,n}$ be whatever it would be (it doesn’t matter because of how we’ll define $\widehat{\mu}^{C}_{-k,n}$ in terms of $\widehat{\mu}_{-k,n}$), and let $P_{k,n}$ consist of $x_{1},\ldots,x_{n/K}$. Recall that $\widehat{\mu}^{C}_{-k,n}$ can depend on $x_{1},\ldots,x_{n/K}$. Here, we define
$\widehat{\mu}^{C}_{-k,n}(x)=\widehat{\mu}_{-k,n}(x)+\sum_{i=1}^{n/K}\mathbf{1}\left\{x=x_{i}\right\}$
so that $\widehat{\mu}^{C}_{-k,n}=\widehat{\mu}_{-k,n}+\sum_{i=1}^{n/K}\delta(x_{i})$ with $\delta$ the dirac delta function. Then
$P|\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X)|=0$
$P_{k,n}|\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X)|=1$
so that
$(P_{k,n}-P)|\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X)|=1$
so that Assumption H does not hold. However, Assumption G still holds as
$P|\widehat{\mu}^{C}_{-k,n}(X)-\widehat{\mu}_{-k,n}(X)|^{2}=0.$
Even though $\widehat{\mu}^{C}_{-k,n}$ is some weird measure 0 modification of $\widehat{\mu}_{-k,n}$, we can also construct other, less weird modifications: for example, instead of adding dirac deltas $\delta(x_{i})$, we could instead add kernels of height 1 around $x_{1},\ldots,x_{n/K}$ and decreasing width (while adjusting kernels to make sure kernels for different $x_{i},x_{j}$ do not overlap) to again satisfy Assumption G but not Assumption H.